### v2

# Optimal Methods for Unknown Piecewise Smooth Problems I: Convex Optimization

Abstract
We introduce an optimal and nearly parameter-free algorithm for minimizing piecewise smooth (PWS) convex functions under the quadratic growth (QG) condition, where the locations and structure of the smooth regions are entirely unknown. Our algorithm, APEX (Accelerated Prox-Level method for Exploring Piecewise Smoothness), is an accelerated bundle-level method designed to adaptively exploit the underlying PWS structure. APEX enjoys optimal theoretical guarantees, achieving a tight oracle complexity bound that matches the lower bound established in this work for convex PWS optimization. Furthermore, APEX generates a verifiable and accurate termination certificate, enabling a robust, almost parameter-free implementation. To the best of our knowledge, APEX is the first algorithm to simultaneously achieve the optimal convergence rate for PWS optimization and provide certificate guarantees.
1 Introduction
Nonsmooth objective functions are central to modern optimization, particularly in machine learning and operations research. Yet, non-smoothness remains a major hurdle for the first-order methods required in large-scale settings; the linear convergence rates of smooth, strongly convex problems [45, 65] deteriorate to sublinear rates for their nonsmooth counterparts. This fundamental gap is well-documented in lower complexity theory [61]. To overcome this limitation, algorithms must exploit the underlying structure of the objective. In this work, we investigate how to utilize piecewise-smooth structures to bridge the divide between smooth and nonsmooth optimization.
Specifically, we consider the problem of minimizing a nonsmooth convex function $f(x)$ over a convex set $X$:
$f^{*}:=\min_{x\in X}f(x).$ (1.1)
Under the assumption that the objective $f$ satisfies the quadratic growth (QG) condition (see Definition 2), we investigate the oracle complexity of first-order methods for finding an $\epsilon$-optimal solution, defined as the theoretical number of (sub)-gradient evaluations required to reach a point $x$ such that $f(x)-f^{*}\leq\epsilon$. We focus on unknown piecewise smooth (PWS) functions: the domain $X$ admits a finite partition into regions on each of which the function is smooth, but the partition is not known a priori (Definition 1). Such functions arise, for example, in quadratic Model Predictive Control (MPC) [42] and two-stage stochastic linear programming (SLP) [16], where those regions are determined by the active constraints and the optimal dual solution in the second-stage problem, respectively, and hence are not available a priori. They also arise in ReLU networks [59], where regions are determined by the composition of activation functions. As network depth increases, these pieces become more difficult to characterize a priori.
The challenge is to design algorithms that automatically exploit this PWS structure to surpass the performance of generic nonsmooth methods [27, 28] to match the performance of smooth optimization. While existing work has predominantly addressed the special case where the objective is a maximum of smooth components [13, 62], the general setting of unknown PWS structure remains largely unexplored. In particular, three major gaps remain in the current literature: the lack of optimal global convergence guarantees, sensitivity to unavailable problem-specific parameters, and the lack of efficient, accurate, and verifiable certificate generation.
Absence of optimal global convergence rates for unknown PWS optimization.
Existing literatures provide several approaches to PWS optimization, frequently leveraging the fact that convex PWS functions can be represented as the maximum of smooth component functions. One line of work [13, 65] directly targets this representation; when gradients for all underlying components are available, prox-linear methods [65, 73, 43, 49] can achieve global linear convergence. In the more general ”unknown” setting, where only the gradient of the currently active component is accessible, the survey descent method by Han and Lewis [33] utilizes gradients at multiple ”survey points” to capture the local max-of-smooth structure. While this approach achieves local linear convergence, its global convergence properties remain missing.
Another research direction with significant empirical success is the bundle method [22, 66, 47, 38, 56]. It aggregates past subgradients to build a cutting-plane model of the objective, and generates new iterates by projecting a candidate point onto a level set defined by this model. Theoretical justification for its superior empirical performance was recently established in the seminal work of Zhang and Sra [74], which proved that the bundle-level (BL) method [47] solves unknown PWS problems with the same oracle complexity as smooth optimization.
However, as an unaccelerated method, the BL method is inherently sub-optimal for smooth problems. It remains unclear whether its accelerated counterpart, the APL method [44], can effectively leverage PWS structures. Our preliminary empirical evidence, illustrated in Figure 2, suggests a negative answer: the theoretically slower BL method consistently outperforms both APL and its restarted variants on the MAXQUAD problem, a prototypical example of PWS functions. This performance discrepancy highlights the need for a new algorithmic framework that bridges the gap between acceleration and structural exploitation. We aim to design an accelerated method that effectively utilizes unknown piecewise structures while achieving optimal convergence rates.
Sensitivity to unavailable problem-specific parameters.
Efficient accelerated algorithms typically require knowledge of the QG modulus $\mu$ to achieve fast convergence. However, because the QG condition is defined globally, $\mu$ is often difficult to estimate in practice [43]. Although Lan et al. recently proposed parameter-free methods for smooth optimization that operate without knowing $\mu$ [43], it remains an open question whether similar guarantees can be achieved in the more challenging PWS setting.
Beyond the QG modulus, the convergence of first-order methods is also governed by the empirical Lipschitz smoothness constant, defined at iteration $k$ as:
$\hat{L}_{k}=\frac{2\left[f(x^{k+1})-f(x^{k})-\left\langle f^{\prime}(x^{k}),x^{k+1}-x^{k}\right\rangle\right]}{\left\|x^{k+1}-x^{k}\right\|^{2}}.$ (1.2)
In smooth optimization, the literature generally distinguishes between two types of complexity guarantees: those based on the worst-case empirical smoothness constant, $L_{\max}=\max_{k}\hat{L}_{k}$, and those based on an averaged smoothness constant $L_{\text{avg}}$, which represents a weighted average of $\{\hat{L}_{k}\}$. While analysis in terms of $L_{\max}$ is often more straightforward [50, 43], bounds based on $L_{\text{avg}}$ are typically tighter and more reflective of superior empirical performance [64]. In PWS optimization, the gap between $L_{\max}$ and $L_{\text{avg}}$ can be even more substantial due to spikes in $\hat{L}_{k}$ that occur when iterates cross into different smooth regions (see Example 1). It is therefore natural to seek a parameter-free algorithm that, without requiring line search or knowledge of $\mu$, attains an optimal complexity bound associated with the average empirical Lipschitz smoothness constant, $L_{\text{avg}}$.
Example 1.1: Average vs. worst-case empirical Lipschitz smoothness constant In the PWS setting, the gradient is continuous within each smooth region but may be discontinuous across their boundaries. This discontinuity can cause the empirical Lipschitz smoothness constant $\hat{L}_{k}$ (1.2) to spike when iterates $x^{k}$ and $x^{k+1}$ belong to different regions. Since the worst-case empirical Lipschitz smoothness constant $L_{\max}$ is determined by these spikes, it can be significantly larger than the average $L_{\text{avg}}$. To illustrate this, consider the iterates generated by applying the Polyak step size to the following one-dimensional objective function with two smooth regions, $\{x\geq 0\}$ and $\{x<0\}$: $\min_{x\in\left[-M,+\infty\right]}f(x;i)=(ix+1)\cdot\mathbf{1}_{\left\{x\geq 0\right\}}+e^{x}\cdot\mathbf{1}_{\left\{x<0\right\}}.$ The objective is parameterized by $i\in\mathbb{N}_{+}$ and $M\in\mathbb{R}$. Starting from an initial point $x^{0}=1/i$, we obtain $L_{\max}=i^{2}e^{-1/i}/2$, while $L_{\text{avg}}\rightarrow e^{-M}$ as the method progresses. Consequently, the ratio $L_{\max}/L_{\text{avg}}$ can be arbitrarily large for certain problem instances. Figure 2 illustrates how these two Lipschitz constants evolve over iterations.
Lack of efficient, accurate, and verifiable certificate generation.
In convex optimization, beyond identifying an accurate candidate solution, it is equally important to provide a termination certificate that verifies the solution’s quality, specifically the optimality gap. Such certificates are essential for early termination and for developing adaptive methods that perform in-situ adjustments to unknown problem parameters. In smooth optimization, the gradient norm is the standard choice for this purpose. However, in nonsmooth optimization, despite significant effort, existing approaches suffer from either computational intractability or limited accuracy. For instance, the gradient norm is no longer a reliable indicator of optimality, the (approximate) Moreau stationarity [58, 18] is expensive to compute, and Goldstein stationarity [31, 17, 52] cannot readily accommodate the set constraint $X$. For PWS optimization, a more recent approach called the Normalized Wolfe Certificate (W-certificate) [74] partially addresses these issues and has been used to design algorithms adaptive to an unknown QG modulus $\mu$. Nevertheless, this certificate remains relatively inaccurate, as the guaranteed optimality gap can be off by a factor of $O(L/\mu)$. This inaccuracy leads to suboptimal oracle complexity bounds for the resulting adaptive methods [74]. This gap in the literature prompts us to ask whether it is possible to efficiently generate a termination certificate that is accurate to within a constant $O(1)$ factor.
Figure 1: Convergence comparison between BL, APL, and Restarted APL on a randomly generated MAXQUAD problem.
Figure 2: Worst and average empirical Lipschitz smoothness constant comparison when running Polyak Step Size in Example 1 with parameter $i=5$, $M=10^{10}$.
In light of these observations, the core challenges of unknown PWS optimization can be distilled into a single, fundamental question:
Can we design a parameter-free algorithm that achieves the optimal convergence rate for unknown PWS optimization while simultaneously generating an accurate termination certificate?
In this work, we answer this question in the affirmative by introducing APEX, a novel BL method that systematically addresses all three major gaps identified above.
Table 1: Various properties comparison between different methods for PWS optimization. $\checkmark$”: can handle unknown PWS optimization; $\dagger$”: using the proof technique developed in our paper, we obtain average empirical Lipschitz smoothness constant dependence for APL; however, the original analysis in [44] yields dependence on the worst-case empirical Lipschitz smoothness constant. QCQP: quadratic constraint quadratic programming; QP: quadratic programming; bounded LP: linear programming within a bounded region.
Methods Unknown Subproblem Oracle
PWS types complexity
Survey Descent [33] $\checkmark$ QCQP local linear rate
BL [74] $\checkmark$ QP $O(1)\left(k^{2}L_{\mathrm{avg}}/\mu^{*}\cdot\log{1}/{\varepsilon}\right)$
APL [44] unknown QP + bounded LP $O(1)\left(\sqrt{L_{\mathrm{avg}}}/{\sqrt{\varepsilon}}\right)^{\dagger}$
APEX (ours) $\checkmark$ QP $O(1)\left(k\sqrt{L_{\mathrm{avg}}/\mu^{*}}\cdot\log{1}/{\varepsilon}\right)$
Our contributions are organized around four main themes:
  1. 1.
    Optimal convergence rate for PWS optimization. We introduce an accelerated bundle-level method termed Apex (Accelerated Prox-Level method for Exploring piecewise-smooth structure) in Section 3. We prove that with an adequate bundle of cutting planes, Apex achieves an optimal iteration complexity of $O\left(\kappa\sqrt{L/\mu^{*}}\log(1/\epsilon)\right)$ to compute an $\epsilon$-optimal solution for objectives satisfying $(\kappa,L)$-PWS (Definition 1) and $\mu^{*}$-QG (Definition 2) conditions. This result significantly improves upon the $O\left(\kappa^{2}L/\mu^{*}\log(1/\epsilon)\right)$ complexity reported in [74]. Our theoretical advancements translate to superior practical performance, as demonstrated in Figure 2. Furthermore, in Section 7, we establish the first lower complexity bound for PWS optimization, defining the theoretical benchmark for optimality. Consequently, Apex is the first algorithm to achieve an accelerated, provably optimal rate for piecewise-smooth optimization.
  2. 2.
    Accelerated certificate generation for PWS optimization.
    In Section 4, we introduce an accelerated bundle-level (BL) framework capable of constructing the W-certificate (see Definition 3) with significantly higher accuracy than previous methods [74]. While prior certification techniques could only bound the optimality gap within a factor of $O(L/\mu)$, our approach improves this bound to a constant factor, $O(1)$. This enhancement allows the certificate to serve as a robust proxy for the true optimality gap. Consequently, the certificate becomes a critical tool for dynamically updating the target level (an estimate of the optimal objective value, $f^{*}$), a essential parameter for the bundle-level method. Leveraging this accurate dynamic update, we design a new parameter-free algorithm, as detailed in our next contribution.
  3. 3.
    Adaptive and almost parameter-free methods for PWS optimization. We develop rAPEX in Section 6, a restarted and almost parameter-free variant of APEX. It only requires that the number of cuts used in the cutting-plane model to exceed the number of pieces in the PWS function. It adaptively exploits the problem structure and provably achieves the optimal complexity with respect to the average empirical Lipschitz smoothness constant, making it both theoretically sound and practically efficient. To our knowledge, rAPEX is the first algorithm to simultaneously be optimal for QG and unknown PWS optimization. A detailed comparison of the properties of different methods for PWS optimization is shown in Table 1.
  4. 4.
    Promising practical performance. We validate our approach through experiments on two representative PWS tasks: MAXQUAD and two-stage stochastic linear programming problems. The MAXQUAD task, a simulated setting, allows precise control over both the smoothness and the number of pieces in the PWS function, enabling rigorous verification of our theoretical claims. For two-stage stochastic linear programming, a widely studied real-world application, we assess the practical performance of our algorithm. Across both tasks, our results demonstrate that APEX delivers strong practical effectiveness for unknown PWS optimization problems.
1.1 Related Work
Algorithms for nonsmooth optimization:
  1. 1)
    Bundle methods. Bundle methods, rooted in Kelley’s cutting-plane scheme [35], iteratively refine a piecewise linear approximation of the objective. Two main variants have been developed: the proximal bundle method and the bundle-level method, distinguished by whether the linear model enters the subproblem as a part of the objectives (proximal) or as a constraint (level).
    1. (a)
      Proximal bundle method. The proximal bundle method was introduced in the 1970s [48, 57] and usually comprises two parts: one is a serious step and one is a null step. Serious steps are used to ensure a decrease of the objective, and null steps are used to refine the information of the descent direction [1]. The complexity analysis for the proximal bundle method has been widely studied in works [39, 29, 26, 51]. For convex functions admitting $\mathcal{VU}$ structure [54], a variant of bundle methods that incorporates second-order information relative to a smooth subspace enjoys superlinear convergence in terms of serious steps [55]. Furthermore, the proximal bundle method has been extended to handle inexact oracle [21] and non-convex objective [34, 23].
    2. (b)
      Bundle level method. Introduced by Lemaréchal, Nesterov, and Nemirovski [47], the bundle-level (BL) method achieves an oracle complexity of $O(1/\epsilon^{2})$ for Lipschitz continuous nonsmooth convex problems. Since then, it has been extended to constrained convex programs, saddle-point problems, and variational inequalities [38]; generalized through non-Euclidean Bregman divergences [5]; accelerated via Nesterov’s techniques [44]; and adapted to function-constrained settings [24]. More recently, Zhang and Sra [74] showed that the BL method attains linear convergence under the QG condition in PWS optimization. However, the resulting complexity’s dependence on the condition number is suboptimal.
  2. 2)
    Goldstein method. The seminal work of Goldstein [31] introduced the $\delta$-Goldstein subdifferential: $\partial_{\delta}f\left(x\right):=\operatornamewithlimits{conv}\left(\cup_{y\in\mathcal{B}(x;\delta)}\partial f(y)\right)$ and defined a $\left(\delta,\varepsilon\right)$-Goldstein stationary point as one satisfying $\min\left\{\left\|g\right\|:g\in\partial_{\delta}f(x)\right\}\leq\varepsilon$. Goldstein [31] originally proposed a subgradient method using the minimal-norm element of $\partial_{\delta}f(x)$, which guarantees descent but is intractable to compute in practice. The gradient sampling algorithm [9, 8, 10] circumvents this by approximating the $\delta$-Goldstein subdifferential through random sampling within $\mathcal{B}(x;\delta)$. More recent work [72, 17, 70] develops efficient randomized approximations for finding a $\left(\delta,\varepsilon\right)$-Goldstein stationary point. Recently, Davis and Jiang [20] further proposed a randomized scheme of the Goldstein method that achieves nearly local linear convergence by exploring the underlying smooth substructures. Kong and Lewis [40] identify abstract properties that enable such near-linear convergence in Goldstein-type methods.
  3. 3)
    Structured nonsmooth optimization. To address the challenges of nonsmoothness, considerable effort has focused on exploiting problem structures. Sparse optimization is a classical example, where composite optimization methods [2, 63] handle objectives combining a smooth term with a prox-friendly nonsmooth regularizer. Problems involving the maximum of smooth functions can be addressed by prox-linear methods [27, 28]. Various smoothing approaches [3, 6, 62, 30, 14] have also been widely applied to mitigate nonsmoothness. Nesterov [62] introduced smoothing techniques for structured nonsmooth problems. Moreau Envelope smoothing [19, 18] and randomized smoothing [30] have received substantial attention. More recently, Li and Cui [49] proposed a subgradient regularization method for nonsmooth marginal functions, achieving notable progress.
Certificate guarantees:
In smooth optimization, the gradient norm is a widely accepted optimality certificate [43]. It offers two essential features. First, it is easily verifiable: given any candidate $\bar{y}$, one can directly compute $\left\|\nabla f(\bar{y})\right\|$ without requiring additional problem parameters. Second, for an $L$-smooth function satisfying the $\mu$-strongly convex property, the gradient norm tightly characterizes the optimality gap:
$\frac{1}{2L}\left\|\nabla f\left(\bar{y}\right)\right\|^{2}\leq f\left(\bar{y}\right)-f^{*}\leq\frac{1}{2\mu}\left\|\nabla f\left(\bar{y}\right)\right\|^{2}\ .$ (1.3)
For nonsmooth but Lipschitz functions, verifiability is preserved, but the strong guarantees in (1.3) no longer hold. By Rademacher’s theorem, Lipschitz functions are differentiable almost everywhere, which motivates defining the subdifferential as $\partial f(x):=\operatornamewithlimits{conv}\left\{g:g=\lim_{t\to\infty}f^{\prime}(x^{t}),x^{t}\to x\right\}$. With this definition, a point $x$ is stationary if $0\in\partial f(x)$. More generally, $x$ is a $(\delta,\varepsilon)$ near-approximate stationary point if $\left\|x-x^{\prime}\right\|\leq\varepsilon$ for some $\varepsilon$-stationary point $x^{\prime}$. While Tian and So [69] introduce the first oracle-polynomial-time algorithm for detecting near-approximate stationary points in piecewise affine functions, for general Lipschitz and bounded-below functions, finding such points can not be guaranteed in polynomial time [41]. As a relaxation, recent work [72, 41] studies $(\delta,\varepsilon)$-Goldstein stationarity, which considers the convex hull of subgradients in a $\delta$-neighborhood. Given a finite sample $\mathcal{P}=\left\{x^{i}\right\}_{i=1}^{m}\subseteq\mathcal{B}(x;\delta)$, one can test $(\delta,\varepsilon)$-Goldstein stationarity by: $\min_{\lambda\in\Delta_{+}}\left\|\sum_{i=1}^{m}\lambda_{i}f^{\prime}(x^{i})\right\|$, where $\Delta_{+}$ is the probability simplex. Approximate Goldstein stationarity is thus computationally tractable, though it does not imply bounds on the optimality gap. On the other hand, for structured problems such as weakly convex functions, the Moreau stationary certificate is widely used. The Moreau envelope [58] is defined as: $f_{\lambda}(x):=\min_{y}\left\{f(y)+\lambda\left\|y-x\right\|^{2}/2\right\}$, which is smooth when $\lambda$ exceeds the weak convexity modulus, with $\nabla f_{\lambda}(x)=\lambda\left(x-x_{\lambda}\right)$ and $x_{\lambda}=\operatornamewithlimits{argmin}_{y}\left\{f(y)+\lambda\left\|y-x\right\|^{2}/2\right\}$. The main challenge is that computing $x_{\lambda}$ exactly is as hard as solving the original problem [19]. Approximating Moreau stationary [6] is possible, but they require an approximation $\hat{x}_{\lambda}$ with $\left\|\hat{x}_{\lambda}-{x}_{\lambda}\right\|$ sufficiently small, which in turn demands an additional termination certificate for the proximal subproblem. Additionally, Nemirovski et al. [60] developed an accuracy certificate framework for convex optimization, variational inequalities with monotone operators, and convex Nash equilibrium problems, showing how such certificates can be constructed via ellipsoid or cutting-plane methods. Nevertheless, this framework has two crucial limitations: it only applies to bounded domains, and it does not accommodate error bounds such as the QG property. As a result, it produces only sublinear convergence rates for certificate generation.
1.2 Outline
The remainder of this paper is organized as follows. In the rest of this section, we introduce the basic notation. Section 2 provides key definitions and background on the bundle-level method. In Section 3, we propose APEX, which drives the objective down to a prescribed level with a convergence rate that adapts to the PWS structure. To remove the requirement of knowing level in advance, Section 4 develops an accelerated Wolfe-certificate search procedure. Building on this certificate, we then study PWS functions satisfying a $\mu^{*}$-QG condition: first assuming $\mu^{*}$ is known (Section 5), and then treating the case where $\mu^{*}$ is unknown (Section 6). In Section 7, we derive a lower bound on the oracle complexity that matches the upper bounds obtained in Sections 5 and 6, thereby establishing optimality of our methods. Finally, Section 8 presents numerical experiments that validate the practical performance of our algorithms.
1.3 Notations
Throughout the paper, we adopt the following notations. Let $n]:=\{1,\ldots,n\}$ for integer $n$. For integer $t$ and $m<t$, let $t-m,t]:=\left\{t-m,t-m+1,\ldots,t\right\}$. The $l_{q}$-norm of a vector $v\in\mathbb{R}^{n}$ is defined as $\left\|v\right\|_{q}=\left(\sum_{i=1}^{n}|v_{(i)}|^{q}\right)^{1/q},$ where $v_{(i)}$ is the $i\text{-th}$ entry of vector $v$. For brevity, $\left\|\cdot\right\|$ denotes the $l_{2}$-norm. The inner product operator for two vectors is defined as $\left\langle u,v\right\rangle=\sum_{i=1}^{n}u_{(i)}v_{(i)}$. We denote the distance between $x$ and a set $X$ as $\operatorname{dist}(x,X)=\min_{\bar{x}\in X}\left\|x-\bar{x}\right\|.$ For brevity, we use $f^{\prime}(x)$ to denote a sub-gradient of $f\left(x\right)$ which coincides with the gradient when $f\left(x\right)$ is differentiable. The linearization of a function $f$ at point $\bar{x}$ is given by $\ell_{f}\left(x;\bar{x}\right):=f\left(\bar{x}\right)+\left\langle f^{\prime}\left(\bar{x}\right),x-\bar{x}\right\rangle.$ The indicator function is defined as $\mathbf{1}_{\left\{A\right\}}={\footnotesize\begin{cases}1,&\text{if }A\text{ is true}\\ 0,&\text{otherwise}\end{cases}}$. The Euclidean ball centered at $\bar{x}$ with radius $\iota$ is denoted as $\mathcal{B}\left(\bar{x},\iota\right)=\left\{x:\left\|x-\bar{x}\right\|\leq\iota\right\}$.
2 Preliminaries
In this section, we first provide several foundational definitions and then introduce the basic BL method for convex nonsmooth optimization. This background will help in understanding the algorithm we propose later.
We begin with the key definitions of $\left(k,L\right)$-piecewise smoothness and $\mu^{*}$-quadratic growth.
Definition 1 ($(k,L)$-Piecewise smoothness, $\left(k,L\right)$-PWS).
We say a function $f:X\to\mathbb{R}$ is $(k,L)$-piecewise smooth if there exists a covering of its domain $X$ by $k$ pieces $\left\{X_{i}\right\}_{i=1}^{k}$ such that for each piece $X_{i}$, the restriction $f_{i}:=f_{\mid X_{i}}$ is $L$-smooth for some $L>0$. Specifically, we assume access to a first order oracle $f^{\prime}(x)$ such that the following inequality holds for all $i\in\left[k\right]:f\left(x\right)-\ell_{f}(x;\bar{x})\leq L\left\|x-\bar{x}\right\|^{2}/2\ ,\ \forall x,\bar{x}\in X_{i}$.
We illustrate two examples of $\left(k,L\right)$-PWS functions in Figures 4 and 4. Figure 4 shows that smooth pieces can be disconnected. Figure 4 provides a simple illustration of the PWS structure in the two-dimensional case. Notably, optimizing such PWS objectives with the Polyak step size may lead to zigzagging behavior [74].
Figure 3: $\max\left\{{x}^{2},x\right\}$ is (2, 2)-PWS and region $X_{1}$ is disconnected.
Figure 4: Trajectory comparison of Polyak step size and BL method on minimizing $\left\|x\right\|^{2}+|x_{(1)}|$ starting from $x^{1}=(0.01,0.15)$.
Definition 2 ($\mu^{*}$-Quadratic growth, $\mu^{*}$-QG).
We say a function $f:X\to\mathbb{R}$ is $\mu^{*}$-quadratic growth if $\mu^{*}=\sup\left\{\mu:f\left(x\right)-f^{*}\geq{\mu}\operatorname{dist}^{2}\left(x,X^{*}\right)/2,\forall x\in X\right\}$ holds, where $X^{*}$ is the nonempty minimizer set, i.e., $X^{*}=\operatornamewithlimits{argmin}_{x\in X}f\left(x\right)$.
We briefly review the BL method, introduced in [47] to improve the stability of the cutting-plane method. At each iteration, the BL method projects the current point onto a specific level set derived from the cutting plane approximation. Specifically, the subproblem of BL is a diagonal quadratic programming problem:
$x^{t}=\operatornamewithlimits{argmin}_{x\in X}\left\|x-x^{t-1}\right\|^{2}\ ,\ \operatornamewithlimits{s.t.}\ \ell_{f}\left(x;x^{i}\right)\leq\tilde{l}\ ,\ i\in[t-m,t-1]\ ,$
where $m$ is the prespecified number of cuts. If we set $\tilde{l}=f^{*}$ and $m=1$, the BL method is equivalent to the Polyak step size [68, 67]. Figure 4 shows that the BL method delivers substantially greater stability than the Polyak method when optimizing PWS problems. This improvement arises because the Polyak step size relies solely on the current gradient, and in PWS objectives with a ”V”-shaped structure, the negative gradient direction often ends up nearly orthogonal to the overall direction of descent, causing pronounced zigzagging. In contrast, the BL method constructs a local model using multiple cutting planes, which more accurately captures the local geometry of the objective and thus supports a much more stable descent direction. Empirically, as demonstrated in Figure 4, the BL method outperforms Polyak step size in terms of trajectory stability for PWS optimization. Theoretically, the superior stability of the BL method over the Polyak step size can be traced to the nature of their convergence analyses. For the Polyak step size, convergence is typically established through a descent property between adjacent iterates (obeying the analysis of BL with $m=1$):
$\begin{aligned} f(x^{t+1})\leq&\ell_{f}(x^{t+1};x^{t})+{\hat{L}_{t}}\left\|x^{t+1}-x^{t}\right\|^{2}/2\\ \leq&f^{*}+{\hat{L}_{t}}(\left\|x^{t}-x^{*}\right\|^{2}-\left\|x^{t+1}-x^{*}\right\|^{2})/2\end{aligned}\ ,$ (2.1)
where $\hat{L}_{t}$ represents the empirical Lipschitz smoothness constant estimated between $x^{t}$ and $x^{t+1}$. Hence, the convergence is strongly dependent on the empirical Lipschitz smoothness constant of adjacent iterates. When two adjacent iterates fall into different smooth pieces, $\hat{L}_{t}$ can be very large due to potential gradient discontinuities, leading to poorer theoretical and practical convergence rates, precisely as reflected in Polyak step size experiments (see Figure 4). In contrast, the BL method permits a refined analysis that leverages non-adjacent iterates. Specifically, its convergence can be established via inequalities of the form [74]:
$\begin{aligned} f(x^{t+i})\leq&\ell_{f}(x^{t+i};x^{t})+{\hat{L}_{t,i}}\left\|x^{t+i}-x^{t}\right\|^{2}/2\\ \leq&f^{*}+i\cdot{\hat{L}_{t,i}}\left(\left\|x^{t}-x^{*}\right\|^{2}-\left\|x^{t+i}-x^{*}\right\|^{2}\right)/2\end{aligned}\ ,\forall i\in\left[m\right]\ ,$ (2.2)
where $\hat{L}_{t,i}$ is the empirical Lipschitz smoothness constant estimated between $x^{t}$ and $x^{t+i}$. Crucially, if the number of cuts exceeds the number of smooth pieces, the relevant Lipschitz dependence is determined by the worst constant among the individual pieces encountered, rather than the potentially much larger empirical Lipschitz smoothness constant across all iterates. This structural advantage explains the improved stability and convergence behavior observed for the BL method in PWS optimization.
Building on the foundation of the BL method, Lan [44] introduced the Accelerated Bundle-Level (ABL) method, which enhances BL by incorporating Nesterov’s acceleration technique. The core iteration includes three steps:
$\begin{split}\underline{x}^{t}&=\left(1-\alpha_{t}\right)\bar{x}^{t-1}+\alpha_{t}x^{t-1}\\ x^{t}&=\operatornamewithlimits{argmin}_{x\in X}\left\|x-x^{t-1}\right\|^{2}\ ,\ \operatornamewithlimits{s.t.}\ \ell_{f}\left(x;\underline{x}^{i}\right)\leq\tilde{l}\ ,\ i\in\left[t\right]\ ,\\ \bar{x}^{t}&=\left(1-\alpha_{t}\right)\bar{x}^{t-1}+\alpha_{t}x^{t}\end{split}$ (ABL)
A distinctive aspect of Nesterov-style acceleration is that the first-order model is evaluated at the weighted-average point $\underline{x}^{t}$, rather than merely at the most recent iterate. This strategy enables the algorithm to effectively incorporate momentum and leads to accelerated convergence rates. ABL is uniformly optimal for smooth, weakly smooth, and nonsmooth functions. However, one critical issue is that, as the algorithm proceeds, the number of cuts $m=t$ increases, which makes the subproblem increasingly difficult to solve. To address this issue, Lan [44] presents APL, which projects from a fixed reference point $\bar{y}$ instead of the current point $x^{t-1}$ and introduces an additional linear constraint (from [4, 5, 38]) to avoid keeping all the cuts. Specifically, it changes the second step of ABL to:
$\begin{split}x^{t}=\operatornamewithlimits{argmin}_{x\in X}\left\|x-\bar{y}\right\|^{2}\ ,\ \operatornamewithlimits{s.t.}\ \left\langle x-x^{t-1},x^{t-1}-\bar{y}\right\rangle\geq 0\ ,\ \ell_{f}\left(x;\underline{x}^{i}\right)\leq\tilde{l}\ ,\ i\in\left[t-m,t-1\right]\ .\end{split}$
This additional linear constraint is introduced to prevent the number of cutting planes from growing too large. A key aspect of the convergence analysis for both ABL and APL lies in their reliance on a local Lipschitz-smoothness condition between adjacent points [44], specifically,
$f(\bar{x}^{t})\leq$ $\ell_{f}(\bar{x}^{t};\underline{x}^{t})+{\hat{L}_{t}}\left\|\bar{x}^{t}-\underline{x}^{t}\right\|^{2}/2$ (2.3)
$\leq$ $\alpha_{t}f^{*}+(1-\alpha_{t})f(\bar{x}^{t-1})+{\hat{L}_{t}}\alpha_{t}^{2}\left(\left\|x^{t}-\bar{y}\right\|^{2}-\left\|x^{t-1}-\bar{y}\right\|^{2}\right)/2$
where $\hat{L}_{t}$ is the empirical Lipschitz smoothness constant determined between $\bar{x}^{t}$ and $\underline{x}^{t}$. A key structural relationship, $\bar{x}^{t}-\underline{x}^{t}=\alpha_{t}(x^{t}-x^{t-1})$, forms the backbone of the telescoping arguments in the convergence proofs of ABL and APL. However, this identity is valid only for adjacent pairs of iterates and fails to generalize to non-adjacent sequences, such as those arising in the BL method discussed in (2.2). This limitation becomes especially pronounced when analyzing PWS functions, where telescoping across non-adjacent points is required.
3 Accelerated Prox-level method for piecewise smooth optimization
We introduce APEX (Accelerated Prox-level method for Exploring Piecewise Smooth), an accelerated algorithm tailored for convex PWS optimization. Building upon the APL framework [44], APEX exploits the PWS structure through a double-loop design: each outer iteration (called a One-Step, see Algorithm 1) comprises $m$ inner steps, where $m$ (the bundle size) determines the number of cutting planes maintained. Similar to APL, Algorithm 1 computes gradients at the averaged points $\underline{x}^{t,i}$ and projects from a fixed reference point $\bar{y}$ onto the level set $X(t,i)$. Additionally, we ensure that the selected solution $\hat{x}^{t}$ forms a non-increasing sequence of objective values across inner iterations. A central distinction from APL is that APEX updates the weighting factor $\alpha_{t}$ only in the outer loop, keeping the averaged point $\hat{x}^{t}$ unchanged within each outer iteration. This approach enables piecewise-adaptive progress towards the target level while retaining the benefits of acceleration. Furthermore, APEX incorporates calculations for certificate generation and empirical Lipschitz estimation, features that are particularly valuable for developing parameter-free algorithms, as discussed in subsequent sections.
Algorithm 1 $\text{One-Step}\left(f,\bar{y},\hat{x}^{t-1},x^{t,0},\tilde{l},m\right)$
1:Initialize: $\text{{Flag}}=\text{{\text{False}}},x^{t+1,0}=x^{t,0}$
2:for $i=1,\ldots,m$ do
3:  $\underline{x}^{t,i}=\left(1-\alpha_{t}\right)\hat{x}^{t-1}+\alpha_{t}x^{t,i-1}$
4:  Set $\underline{X}(t,i)=\left\{x\in X:\ell_{f}\left(x;\underline{x}^{t,i}\right)\leq\tilde{l}\right\}$ and $\bar{X}(t,i-1)=\left\{x:\left\langle x-x^{t,i-1},x^{t,i-1}-\bar{y}\right\rangle\geq 0\right\}$
5:  Set $X(t,i):=\left\{\cap_{j\in\left[1,i\right]}\underline{X}(t,j)\right\}\cap\bar{X}\left(t,0\right)$
6:  Compute
$x^{t,i}\leftarrow\operatornamewithlimits{argmin}_{x\in X(t,i)}\left\|x-\bar{y}\right\|^{2}$ (3.1)
7:  if the subproblem is infeasible then
8:   Set $\text{{Flag}}=\text{{True}}$ and go to Line 13
9:  end if
10:  Compute $\bar{x}^{t,i}=\left(1-\alpha_{t}\right)\hat{x}^{t-1}+\alpha_{t}x^{t,i}$ and set point set $\mathcal{P}=\left\{\bar{y},\underline{x}^{t,1},\ldots,\underline{x}^{t,i}\right\}$
11:  Choose $\hat{x}^{t}$ such that $f\left(\hat{x}^{t}\right)=\min\left\{f\left(\hat{x}^{t-1}\right),\min_{j\leq i}\left\{f\left(\bar{x}^{t,j}\right)\right\}\right\}$ and $x^{t+1,0}=x^{t,i}$
12:end for
13: Compute:
$\left(r_{t},l_{t}\right)=\operatornamewithlimits{argmin}_{0\leq l<r\leq m}L_{t}\left(r,l\right),\text{where}\ L_{t}\left(r,l\right)=\left(f(\hat{x}^{t})-\tilde{l}-(1-\frac{3}{4}\alpha_{t})(f(\hat{x}^{t-1})-\tilde{l})\right)\big/\left(\frac{\alpha_{t}^{2}}{2}\|x^{t,r}-x^{t,l}\|^{2}\right)$
14:Return: $\hat{x}^{t},x^{t+1,0},\text{{Flag}},\mathcal{P},L_{t}\left(r_{t},l_{t}\right),\left\|x^{t,r_{t}}-x^{t,l_{t}}\right\|$
Algorithm 2 Accelerated Prox-level Method for Exploring Piecewise Smooth (APEX)
1:objective function $f$, the candidate solution $\bar{y}$, test level $\tilde{l}$, and the number of cuts $m$
2:Initialize: $x^{1,0}=\bar{y},\hat{x}^{0}=\bar{y}$
3:for $t=1,\ldots,N$ do
4:  $\left(\hat{x}^{t},x^{t+1},\ldots\right)\leftarrow\text{One-Step}\left(f,\bar{y},\hat{x}^{t-1},x^{t},\tilde{l},m\right)$
5:end for
In contrast to existing analyses of APL [44], which depend on the Lipschitz continuity of the objective function across successive iterates, our proposed mechanism allows the derivation of optimality conditions across non-adjacent iterates. This flexibility enables the convergence proof to bypass certain intermediate points, leading to tighter theoretical guarantees with improved dependence on the Lipschitz constant. We first define
$L_{t}=\min_{0\leq l<r\leq m}L_{t}(r,l)\ \ \text{and}\ \ \bar{L}\left(N\right)=\frac{\sum_{t=1}^{N}\omega_{t}\alpha_{t}^{2}L_{t}\mathbf{1}_{\mathcal{E}(t)}\left\|x^{t,r_{t}}-x^{t,l_{t}}\right\|^{2}/2}{\sum_{t=1}^{N}\mathbf{1}_{\mathcal{E}(t)}\left\|x^{t,r_{t}}-x^{t,l_{t}}\right\|^{2}}\ ,$ (3.2)
[…]
Proposition 3.1.
Suppose function $f$ is convex, $\left(k,L\right)$-PWS and the sequence $\left\{\hat{x}^{t}\right\}$ generated by Algorithm 2 satisfies $f\left(\hat{x}^{t}\right)\geq\tilde{l}$ for all $t\geq 1$ and that every subproblem (3.1) of Algorithm 2 is feasible. Let $x^{t}:=x^{t,m}$, and assume the weight $\left\{\omega_{t}\right\}$ satisfy $\omega_{t}\left(1-\alpha_{t}/2\right)=\omega_{t-1}$ for $t\geq 1$. Then the following properties hold:
  1. 1.
    Monotonicity of Proximal Distances: For every $t\geq 0$ and $0\leq i\leq m$, it holds that $\|x^{t,i}-\bar{y}\|\leq\|x^{t,i+1}-\bar{y}\|$.
  2. 2.
    Convergence Bound: There exists a constant $\bar{L}(t)$ (defined in (3.2)) such that
    $\omega_{t}\left[f(\hat{x}^{t})-\tilde{l}\right]-\omega_{1}\left(1-\frac{\alpha_{1}}{2}\right)\left[f(\hat{x}^{0})-\tilde{l}\right]\leq\bar{L}(t)\|x^{t}-\bar{y}\|^{2}\ .$ (3.3)
Proof.
Part 1 (Monotonicity). We first establish that the distances $\left\|x^{t,i}-\bar{y}\right\|$ are nondecreasing across the inner-loop updates. For each $1\leq i\leq m$, recall that $x^{t,i}\in\operatornamewithlimits{argmin}_{x\in X(t,i)}\left\|x-\bar{y}\right\|^{2}$. By the three-point property for Euclidean projections, this implies
$\left\|x^{t,i}-\bar{y}\right\|^{2}+\left\|x^{t,i}-x\right\|^{2}\leq\left\|x-\bar{y}\right\|^{2}\ ,\ \forall x\in X(t,i)\ .$
Because $x^{t,i+1}\in X(t,i+1)\subseteq X(t,i)$ and $x^{t,i+1}\in X\left(t,i+1\right)$, substituting $x=x^{t,i+1}$ gives
$\left\|x^{t,i}-\bar{y}\right\|^{2}+\left\|x^{t,i}-x^{t,i+1}\right\|^{2}\leq\left\|x^{t,i+1}-\bar{y}\right\|^{2}\ ,$
which implies $\left\|x^{t,i}-\bar{y}\right\|\leq\left\|x^{t,i+1}-\bar{y}\right\|$.
For the initial case $i=0$, we expand
$\begin{split}\left\|x^{t,1}-\bar{y}\right\|^{2}&=\left\|x^{t,1}-x^{t,0}\right\|^{2}+2\left\langle x^{t,1}-x^{t,0},x^{t,0}-\bar{y}\right\rangle+\left\|x^{t,0}-\bar{y}\right\|^{2}\end{split}\ .$
The linear constraint in $X(t,0)$ ensures the cross term is nonnegative, so $\left\|x^{t,0}-\bar{y}\right\|^{2}\leq\left\|x^{t,1}-\bar{y}\right\|^{2}$. Hence $\left\|x^{t,i}-\bar{y}\right\|\leq\left\|x^{t,i+1}-\bar{y}\right\|$, which completes the proof of monotonicity.
Part 2 (Convergence bound). We now turn to the function-value recursion.
Let the index $\left\{\left(l_{1},r_{1}\right),\ldots,\left(l_{t},r_{t}\right)\right\}$ be the minimizer index pair for $\left\{L_{1},\ldots,L_{t}\right\},$where $l_{i}<r_{i},\forall i\in[t].$ By the optimality condition of updating rule for $x^{t,l_{t}}$, we have
$\left\|x^{t,l_{t}}-\bar{y}\right\|^{2}+\left\|x^{t,l_{t}}-x\right\|^{2}\leq\left\|x-\bar{y}\right\|^{2}\ ,\ \forall x\in X\left(t,l_{t}\right)\ .$
Since condition $X(t,r_{t})\subseteq X(t,l_{t})$ holds by the updating rule of $X\left(t,i\right)$. Hence, taking $x=x^{t,r_{t}}\in X(t,r_{t})\subseteq X(t,l_{t})$ in above gives:
$\left\|x^{t,l_{t}}-\bar{y}\right\|^{2}+\left\|x^{t,l_{t}}-x^{t,r_{t}}\right\|^{2}\leq\left\|x^{t,r_{t}}-\bar{y}\right\|^{2}\ .$ (3.4)
Next, we give the descent analysis of the sequence based on (3.4). It follows from the non-decreasing property in Part1 and (3.4) that
$\left\|x^{t,0}-\bar{y}\right\|^{2}+\left\|x^{t,l_{t}}-x^{t,r_{t}}\right\|^{2}\leq\left\|x^{t,m}-\bar{y}\right\|^{2}\ .$
Summing from $t=1$ to $N$ gives
$\sum_{t=1}^{N}\left\|x^{t,l_{t}}-x^{t,r_{t}}\right\|^{2}\overset{(a)}{=}\left\|x^{1,0}-\bar{y}\right\|^{2}+\sum_{t=1}^{N}\left\|x^{t,l_{t}}-x^{t,r_{t}}\right\|^{2}\leq\left\|x^{N,m}-\bar{y}\right\|^{2}\ ,$ (3.5)
where $(a)$ holds by that $x^{1,0}=\bar{y}$. From the definition of $L_{t}$, we obtain
$\begin{split}f\left(\hat{x}^{t}\right)-\tilde{l}&\leq\left(1-\frac{3}{4}\alpha_{t}\right)\left(f\left(\hat{x}^{t-1}\right)-\tilde{l}\right)+\frac{L_{t}\alpha_{t}^{2}}{2}\left\|x^{t,r_{t}}-x^{t,l_{t}}\right\|^{2}\\ &\leq\left(1-\frac{\alpha_{t}}{2}\right)\left(f\left(\hat{x}^{t-1}\right)-\tilde{l}\right)+\frac{L_{t}\alpha_{t}^{2}}{2}\left\|x^{t,r_{t}}-x^{t,l_{t}}\right\|^{2}\end{split}\ .$
Then combining the above inequality and the definition of $\mathbf{1}_{\mathcal{E}(t)}$ gives
$f\left(\hat{x}^{t}\right)-\tilde{l}\leq\left(1-\frac{\alpha_{t}}{2}\right)\left(f\left(\hat{x}^{t-1}\right)-\tilde{l}\right)+\frac{L_{t}\mathbf{1}_{\mathcal{E}(t)}\alpha_{t}^{2}}{2}\left\|x^{t,r_{t}}-x^{t,l_{t}}\right\|^{2}.$
Multiplying $\omega_{t}$ on both sides of it and using $\omega_{t}\left(1-\frac{\alpha_{t}}{2}\right)=\omega_{t-1}$, we obtain
$\begin{split}\omega_{t}\left(f\left(\hat{x}^{t}\right)-\tilde{l}\right)&\leq\omega_{t}\left(1-\frac{\alpha_{t}}{2}\right)\left(f\left(\hat{x}^{t-1}\right)-\tilde{l}\right)+\frac{\omega_{t}L_{t}\mathbf{1}_{\mathcal{E}(t)}\alpha_{t}^{2}}{2}\left\|x^{t,r_{t}}-x^{t,l_{t}}\right\|^{2}\\ &\overset{(a)}{=}\omega_{t-1}\left(f\left(\hat{x}^{t-1}\right)-\tilde{l}\right)+\frac{\omega_{t}L_{t}\mathbf{1}_{\mathcal{E}(t)}\alpha_{t}^{2}}{2}\left\|x^{t,r_{t}}-x^{t,l_{t}}\right\|^{2}\ ,\\ \omega_{1}\left(f\left(\hat{x}^{1}\right)-\tilde{l}\right)&\leq\omega_{1}\left(1-\frac{\alpha_{1}}{2}\right)\left(f\left(\hat{x}^{0}\right)-\tilde{l}\right)+\frac{\omega_{1}L_{1}\mathbf{1}_{\mathcal{E}(1)}\alpha_{1}^{2}}{2}\left\|x^{1,r_{1}}-x^{1,l_{1}}\right\|^{2}\ ,\end{split}$ (3.6)
where $(a)$ holds by the definition of $\omega_{t}$. Summing (3.6) from $t=1$ to $N$ yields
$\omega_{N}\left(f\left(\hat{x}^{N}\right)-\tilde{l}\right)\leq\omega_{1}\left(1-\frac{\alpha_{1}}{2}\right)\left(f\left(\hat{x}^{0}\right)-\tilde{l}\right)+\bar{L}\left(N\right)\sum_{t=1}^{N}\mathbf{1}_{\mathcal{E}(t)}\left\|x^{t,r_{t}}-x^{t,l_{t}}\right\|^{2}\ .$ (3.7)
Combining $\sum_{t=1}^{N}\mathbf{1}_{\mathcal{E}(t)}\left\|x^{t,r_{t}}-x^{t,l_{t}}\right\|^{2}\leq\sum_{t=1}^{N}\left\|x^{t,r_{t}}-x^{t,l_{t}}\right\|^{2}$ and (3.7) gives
[…]
where $(a)$ is by (3.5). Hence, we complete our proof.
We make a few remarks regarding the result.
First, Algorithm 2 is a double-loop method. By updating the weighting parameter $\alpha_{t}$ and the reference iterate $\hat{x}^{t}$ only once per outer iteration, the convergence analysis can compare arbitrary pairs of inner-loop iterates, rather than only successive ones, by applying the three-point lemma (see (3.4)). This flexibility enables the algorithm to exploit regions with smaller local Lipschitz constants, thereby better adapting to the PWS.
Second, within each inner iteration, a cut-generation scheme is defined as
$\text{Option I}:\ \ X(t,i):=\left\{\cap_{j\in\left[1,i\right]}\underline{X}(t,j)\right\}\cap\bar{X}\left(t,0\right)\ .$ (3.8)
In this scheme, the algorithm starts with two initial cuts; at each of the $m$ inner steps, it appends one new cutting plane derived from the updated point $\underline{x}^{t,i}$. A major limitation of this strategy is that all cuts from the previous iteration must be discarded when entering a new outer iteration. Alternatively, one may consider
$\text{Option II}:\ \ X\left(t,i\right):=\left\{\cap_{j\in\left[1,i\right]}\underline{X}(t,j)\right\}\cap\bar{X}\left(t-1,0\right)\cap\left\{\cap_{j\in\left[1,m\right]}\underline{X}(t-1,j)\right\}$ (3.9)
Here, we replace $\bar{X}\left(t,0\right)$ with $\bar{X}\left(t-1,0\right)\cap\left\{\cap_{j\in\left[1,m\right]}\underline{X}(t-1,j)\right\}$. This modification preserves all properties established in Proposition 3.1 and does not affect the convergence result. The advantage is that cuts from the previous outer iteration are retained. However, if the number of cuts is fixed, the linear constraint $\left\{\cap_{j\in\left[1,i\right]}\underline{X}(t,j)\right\}$ will be smaller than that in the Option I scheme. Consequently, the Option II approach may reduce the algorithm’s ability to capture the piecewise smooth structure.
Third, Proposition 3.1 contains two important properties. The first is that the $\left\{\left\|x^{t,i}-\bar{y}\right\|^{2}\right\}$ is non-decreasing, which helps in constructing a lower bound of $f^{*}$ in the following Sections, the second property implies that if the domain $X$ is bounded, i.e., there exists a constant $D_{X}>0$ such that $D_{X}=\max_{x_{1},x_{2}\in X}\left\|x_{1}-x_{2}\right\|$, then we have
$\omega_{N}\left(f\left(\hat{x}^{N}\right)-\tilde{l}\right)-\omega_{1}\left(1-\frac{\alpha_{1}}{2}\right)\left(f\left(\hat{x}^{1}\right)-\tilde{l}\right)\leq\bar{L}\left(N\right)D_{X}^{2}\ ,$ (3.10)
which implies $f\left(\hat{x}^{N}\right)-\tilde{l}=\mathcal{O}\left(\frac{\bar{L}\left(N\right)D_{X}^{2}}{\omega_{N}}\right)$ and this is also the core idea behind Lan’s proof in [44].
Fourth, in many applications one may choose the target level $\tilde{l}$ to be the true optimum $f^{*}$; this is exactly the strategy underlying Polyak step-size [68, 67] and Polyak Minorant Method [25]. Furthermore, Dang et al. [15] has applied this idea to semidefinite programming.
Substituting $\tilde{l}=f^{*}$ into the bound of Proposition 3.1, we give the following sub-linear convergence bound.
Theorem 3.1.
Suppose function $f$ is convex and $\left(k,L\right)$-PWS. Let the weighted sequence $\omega_{t}=\left(t+2\right)\left(t+3\right)/2$ and $\alpha_{t}=4/\left(t+3\right)$ in APEX. Then for the sequence $\left\{\hat{x}^{t}\right\}$ generated by APEX with level $\tilde{l}=f^{*}$, there exists a constant $\bar{L}\left(t\right)$ such that
$f\left(\hat{x}^{t}\right)-f^{*}\leq\frac{6}{\left(t+2\right)\left(t+3\right)}\left(f\left(\hat{x}^{0}\right)-f^{*}\right)+\frac{2\bar{L}\left(t\right)}{\left(t+2\right)\left(t+3\right)}\left\|x^{*}-\bar{y}\right\|^{2}\ .$
Proof.
The choice of $\omega_{t}$ and $\alpha_{t}$ satisfy $\omega_{t}\left(1-\frac{\alpha_{t}}{2}\right)=\omega_{t-1}$. Hence, it follows from the second result of Proposition 3.1 that
$\omega_{t}\left(f\left(\hat{x}^{t}\right)-f^{*}\right)-\omega_{1}\left(1-\frac{\alpha_{1}}{2}\right)\left(f\left(\hat{x}^{0}\right)-f^{*}\right)\leq\bar{L}\left(t\right)\left\|x^{t}-\bar{y}\right\|^{2}\ .$
Note that the optimal solution $x^{*}$ is always feasible for the subproblem if we set the level $\tilde{l}=f^{*}$. Hence, any feasible solution to the subproblem $x^{t}$ should satisfy $\left\|x^{t}-\bar{y}\right\|\leq\left\|x^{*}-\bar{y}\right\|$. ∎
One remaining question is that, even though $\bar{L}\left(t\right)$ (the average of $L_{t}$, defined in (3.2)) is computable in practice, it is still unclear how to bound the constant $\bar{L}\left(t\right)$ as it appears in Proposition 3.1 and Theorem 3.1. In the following Proposition 3.2, we show that value $\bar{L}\left(t\right)$ can be upper bounded by a weighted average of empirical Lipschitz smoothness constants $\tilde{L}_{t}$, where $\tilde{L}_{t}$ is the best Lipschitz constant between $\bar{x}^{t,r}$ and $\underline{x}^{t,l+1}$ with $r>l$ in $t$-th outer iteration.
Proposition 3.2.
Let $\left\{\hat{x}^{t},x^{t,i}\right\}$ be the iterates generated by Algorithm 2, and assume that for each outer iteration $t\geq 1$, there exist indices $0\leq l_{t}<r_{t}\leq m$ and a constant $\tilde{L}_{t}>0$ such that
[…]
then $L_{t}\leq\tilde{L}_{t}$ ($L_{t}$ is defined in (3.2)). Furthermore, let the weighted sequence $\omega_{t}$ satisfy $\omega_{t}=\omega_{t-1}\left(1-{\alpha_{t}}/{2}\right)^{-1}$. Then $\bar{L}\left(N\right)$ (defined in (3.2)) admits the upper bound:
$\bar{L}\left(N\right)<\frac{3\sum_{t=1}^{N}\mathbf{1}_{\mathcal{E}(t)}\omega_{t}\left(\alpha_{t}\left(f\left(\hat{x}^{t-1}\right)-\tilde{l}\right)\right)}{\sum_{t=1}^{N}\mathbf{1}_{\mathcal{E}(t)}\left(f\left(\hat{x}^{t-1}\right)-\tilde{l}\right)\Big/\left(\alpha_{t}\cdot\tilde{L}_{t}\right)}\ .$ (3.12)
Proof.
It follows from $f\left(\hat{x}^{t}\right)\leq f\left(\bar{x}^{t,r_{t}}\right)$ and the definition of $\bar{x}^{t,r_{t}}=\left(1-\alpha_{t}\right)\hat{x}^{t-1}+\alpha_{t}x^{t,r_{t}}$ that
[…]
where $(a)$ holds by the definition of $\tilde{L}_{t}$, $(b)$ is by the definition of $\bar{x}^{t,r_{t}}$ and $\ell_{f}$, and $(c)$ follows from the $f$ is convex and $x^{t,r_{t}}\in X\left(t,r_{t}\right)$. Subtracting $\tilde{l}$ on both sides gives
$f\left(\hat{x}^{t}\right)-\tilde{l}\leq\left(1-\frac{3}{4}\alpha_{t}\right)\left(f\left(\hat{x}^{t-1}\right)-\tilde{l}\right)+\frac{\tilde{L}_{t}\alpha_{t}^{2}}{2}\left\|x^{t,r_{t}}-x^{t,l_{t}}\right\|^{2}\ ,$
which implies that $L_{t}\leq\tilde{L}_{t}$. Next, we consider the upper bound of $\bar{L}\left(N\right)$. By the definition of $\bar{L}\left(N\right)$ and $\mathbf{1}_{\mathcal{E}(t)}$ , the upper bound of $\bar{L}\left(N\right)$ is given as follows:
$\begin{split}\bar{L}\left(N\right)&=\frac{\sum_{t=1}^{N}\omega_{t}\alpha_{t}^{2}L_{t}\mathbf{1}_{\mathcal{E}(t)}\left\|x^{t,r_{t}}-x^{t,l_{t}}\right\|^{2}/2}{\sum_{t=1}^{N}\mathbf{1}_{\mathcal{E}(t)}\left\|x^{t,r_{t}}-x^{t,l_{t}}\right\|^{2}}\\ &\overset{(a)}{=}\frac{\frac{1}{2}\sum_{t=1}^{N}\mathbf{1}_{\mathcal{E}(t)}\omega_{t}\left(\left(f\left(\hat{x}^{t}\right)-\tilde{l}\right)-\left(1-\frac{3}{4}\alpha_{t}\right)\left(f\left(\hat{x}^{t-1}\right)-\tilde{l}\right)\right)}{\sum_{t=1}^{N}\mathbf{1}_{\mathcal{E}(t)}\left(\left(f\left(\hat{x}^{t}\right)-\tilde{l}\right)-\left(1-\frac{3}{4}\alpha_{t}\right)\left(f\left(\hat{x}^{t-1}\right)-\tilde{l}\right)\right)\Big/\left(\alpha_{t}^{2}\cdot L_{t}\right)}\\ &\overset{(b)}{<}\frac{\frac{3}{8}\sum_{t=1}^{N}\mathbf{1}_{\mathcal{E}(t)}\omega_{t}\left(\alpha_{t}\left(f\left(\hat{x}^{t-1}\right)-\tilde{l}\right)\right)}{\frac{1}{4}\sum_{t=1}^{N}\mathbf{1}_{\mathcal{E}(t)}\left(f\left(\hat{x}^{t-1}\right)-\tilde{l}\right)\Big/\left(\alpha_{t}\cdot L_{t}\right)}=\frac{\frac{3}{2}\sum_{t=1}^{N}\mathbf{1}_{\mathcal{E}(t)}\omega_{t}\left(\alpha_{t}\left(f\left(\hat{x}^{t-1}\right)-\tilde{l}\right)\right)}{\sum_{t=1}^{N}\mathbf{1}_{\mathcal{E}(t)}\left(f\left(\hat{x}^{t-1}\right)-\tilde{l}\right)\Big/\left(\alpha_{t}\cdot L_{t}\right)}\end{split}\ ,$
where $(a)$ holds by the definition of $L_{t}$ and $(b)$ follows from $f\left(\hat{x}^{t}\right)\leq f\left(\hat{x}^{t-1}\right)$ and the definition of $\mathcal{E}(t)$. Finally, since $L_{t}\leq\tilde{L}_{t}$, the stated upper bound follows. ∎
We make a few remarks regarding the result.
[…]
Second, if the objective $f$ is globally $L$ smooth, then each local surrogate $\tilde{L}_{t}$ is upper bounded by $L$ uniformly, and hence $\bar{L}\left(N\right)=O(1)L$. More generally, when $f$ is $\left(k,L\right)$-PWS with at most $k$ smooth segments and one maintains $m\geq k$ cuts per outer iteration, one again obtains $\bar{L}\left(N\right)=O(1)L$ throughout. However, if $m<k$, then we have the following upper bound by $M$-Lipschitz continuous:
$\begin{split}\tilde{L}_{t}=&2\left(f\left(\bar{x}^{t,r_{t}}\right)-\ell_{f}\left(\bar{x}^{t,r_{t}},\underline{x}^{t,l_{t}+1}\right)-\frac{\alpha_{t}}{4}\left(f\left(\hat{x}^{t-1}\right)-\tilde{l}\right)\right)\Big/\left\|\bar{x}^{t,r_{t}}-\underline{x}^{t,l_{t}+1}\right\|^{2}\\ \leq&2\left(M\left\|\bar{x}^{t,r_{t}}-\underline{x}^{t,l_{t}+1}\right\|-\frac{\alpha_{t}}{4}\left(f\left(\hat{x}^{t-1}\right)-\tilde{l}\right)\right)\Big/\left\|\bar{x}^{t,r_{t}}-\underline{x}^{t,l_{t}+1}\right\|^{2}\\ \overset{(a)}{\leq}&2\left(\frac{M^{2}\left\|\bar{x}^{t,r_{t}}-\underline{x}^{t,l_{t}+1}\right\|^{2}}{\alpha_{t}\left(f\left(\hat{x}^{t-1}\right)-\tilde{l}\right)}\right)\Big/\left\|\bar{x}^{t,r_{t}}-\underline{x}^{t,l_{t}+1}\right\|^{2}=\frac{2M^{2}}{\alpha_{t}\left(f\left(\hat{x}^{t-1}\right)-\tilde{l}\right)}\end{split}\$,
[…]
4 Normalized Wolfe certificate
To guarantee that the iterates $\left\{f\left(\hat{x}^{t}\right)\right\}$ converge to the true optimum $f^{*}$, Algorithm 2 requires prior knowledge of $\tilde{l}=f^{*}$. While this assumption holds in certain applications [15, 25], it is generally impractical in broader optimization settings. To address this limitation, we instead estimate a lower bound of $f^{*}$. Since the current function value provides a natural upper bound, we may set the target level as the midpoint between the lower and upper bounds, thereby approximating $f^{*}$. As optimization progresses, this estimate can be refined, ultimately leading to the exact optimum.
In this section, we introduce the Normalized Wolfe Certificate (W-certificate), a procedure that constructs a reliable lower bound on $f^{*}$ by exploiting the QG condition of the objective. Specifically, we (i) formally define the W-certificate, (ii) prove that it yields a valid underestimate of $f^{*}$ under QG condition, and (iii) describe how to apply Algorithm 3 to generate a W-certificate.
Definition 3 (Normalized Wolfe certificate, W-certificate).
Let $\iota>0$ and $\nu\geq 0$. $\bar{y}\in X$ is called an $\left(\iota,\nu\right)$-Normalized Wolfe stationary of function $f$ if there exists an evaluation point set $\mathcal{P}=\left\{z^{0}=\bar{y},z^{1},\ldots,z^{m}\right\}\subseteq\mathcal{B}\left(\bar{y},\iota\right)$ such that
$\mathcal{V}_{\mathcal{P},f}\left(\iota;\bar{y}\right)=\frac{1}{\iota}\max_{x\in\mathcal{B}\left(\bar{y},\iota\right)\cap X}\left(\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(x\right)\right)\leq\nu\ ,\ \text{where}\ \psi_{\mathcal{P},f}\left(x\right)=\max_{z\in\mathcal{P}}\left\{\ell_{f}\left(x,z\right)\right\}\ .$ (4.1)
In this case, we call $\mathcal{P}$ an $\left(\iota,\nu\right)$-W-certificate for $f$ at point $\bar{y}$.
W-certificate is firstly introduced by [74], the W-certificate extends the classical gradient-norm criterion to PWS objectives by quantifying the maximal normalized decrease of a piecewise-linear surrogate model. This model is constructed from multiple supporting hyperplanes evaluated within an $\iota$-neighborhood of the current iterate, and the resulting certificate can be computed using only a first-order oracle.
Similar to the gradient norm, the W-certificate exhibits a comparable property: 1. W-certificate serves as a practical termination criterion when $f$ is PWS, playing the same role as the gradient-norm in smooth optimization. 2. W-certificate provides a trustworthy lower bound on the optimum $f^{*}$ under the QG condition, which guarantees an upper bound on the true optimality gap, thereby yielding a parameter-free and practically verifiable termination criterion. In what follows, we first establish one fundamental properties of the normalized Wolfe-Gap $\mathcal{V}_{\mathcal{P},f}\left(\iota;\bar{y}\right)$ in Lemma 4.1 , and subsequently demonstrate in Lemma 4.2 that under the QG condition, W-certificate is provably proportional to the true optimality gap.
Lemma 4.1.
$\mathcal{V}_{\mathcal{P},f}\left(\iota;\bar{y}\right)$ (defined in (4.1)) is non-increasing with respect to $\iota$.
Proof.
Let some $\iota_{1}>\iota_{2}$ be given and let $x_{\iota 1}$ and $x_{\iota 2}$ denote the respective optimal solution of $\mathcal{V}_{\mathcal{P},f}\left(\iota_{1};\bar{y}\right)$ and $\mathcal{V}_{\mathcal{P},f}\left(\iota_{2};\bar{y}\right)$. By the convexity of $X$, we have point $\tilde{x}:=\bar{y}+\frac{\iota_{2}}{\iota_{1}}\left(x_{\iota 1}-\bar{y}\right)\in X.$ Since $\psi_{\mathcal{P},f}\left(x\right)$ is convex and $\bar{y},\tilde{x}$ and $x_{\iota 1}$ lie on the same line, we get from the monotonicity of the secant line:
$\frac{\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(\tilde{x}\right)}{\left\|\tilde{x}-\bar{y}\right\|}\geq\frac{\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(x_{\iota 1}\right)}{\left\|x_{\iota 1}-\bar{y}\right\|}\ .$ (4.2)
Combining (4.2), $\left\|\bar{y}-\tilde{x}\right\|=\frac{\iota_{2}}{\iota_{1}}\left\|x_{\iota 1}-\bar{y}\right\|$ gives $\frac{\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(\tilde{x}\right)}{\iota_{2}}\geq\frac{\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(x_{\iota 1}\right)}{\iota_{1}}$. By the definition of $x_{\iota 2}$, we have $\frac{\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(x_{\iota 1}\right)}{\iota_{1}}\leq\frac{\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(\tilde{x}\right)}{\iota_{2}}\leq\frac{\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(x_{\iota 2}\right)}{\iota_{2}}$. Hence, we complete our proof of that $\mathcal{V}_{\mathcal{P},f}\left(\iota;\bar{y}\right)$ is non-increasing with respect to $\iota$. ∎
Lemma 4.2.
Let $f$ be a convex function satisfying $\mu$-QG. If $\mathcal{P}$ constitutes an $\left(\iota,\nu\right)$-W-certificate for the point $\bar{y}$, then the function value at $\bar{y}$ obeys the bound $f\left(\bar{y}\right)-f^{*}\leq\max\left\{\iota\nu,{2\nu^{2}}/{\mu}\right\}$.
Proof.
Since the objective function $f$ is convex, then we have $\psi_{\mathcal{P},f}\left(x\right)\leq f\left(x\right)$. Furthermore, condition $\bar{y}\in\mathcal{P}$ implies $\psi_{\mathcal{P},f}\left(\bar{y}\right)\leq f\left(\bar{y}\right)$. We consider two cases based on the distance between $x_{p}^{*}$ and $\bar{y}$, where $x_{p}^{*}=\operatornamewithlimits{argmin}_{x\in X^{*}}\left\|\bar{y}-x\right\|$ and $X^{*}:=\left\{x\in X:f\left(x\right)=f^{*}\right\}$.
Case 1. $x_{p}^{*}\in\mathcal{B}\left(\bar{y},\iota\right)$. Then we have
$\begin{split}f\left(\bar{y}\right)-f^{*}\leq f\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(x_{p}^{*}\right)&\overset{(a)}{\leq}\max_{x\in\mathcal{B}\left(\bar{y},\iota\right)\cap X}\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(x\right)\end{split}\leq\iota\nu\ ,$
where $(a)$ follows from function $f$ is convex and $\bar{y}\in\mathcal{P}$ and these facts implies that $f\left(\bar{y}\right)=\psi_{\mathcal{P},f}\left(\bar{y}\right)$.
Case 2. $x_{p}^{*}\in\mathcal{B}^{c}\left(\bar{y},\iota\right)$. From the radius argument that $\left\|x_{p}^{*}-\bar{y}\right\|\geq\iota$ and $\mathcal{V}_{\mathcal{P},f}\left(\iota;\bar{y}\right)$ is non-increasing with respect to $\iota$, we have
$\begin{split}f\left(\bar{y}\right)-f^{*}&\leq\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(x_{p}^{*}\right)\\ &\leq\max_{x\in\mathcal{B}\left(\bar{y},\left\|x_{p}^{*}-\bar{y}\right\|\right)}\left\{\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(x\right)\right\}\\ &=\left\|x_{p}^{*}-\bar{y}\right\|\mathcal{V}_{\mathcal{P},f}\left(\left\|x_{p}^{*}-\bar{y}\right\|;\bar{y}\right)\\ &\leq\left\|x_{p}^{*}-\bar{y}\right\|\mathcal{V}_{\mathcal{P},f}\left(\iota;\bar{y}\right)\leq\left\|x_{p}^{*}-\bar{y}\right\|\nu\end{split}\ .$ (4.3)
Combining the above inequality and $f\left(\bar{y}\right)-f^{*}\geq\frac{\mu}{2}\left\|\bar{y}-x_{p}^{*}\right\|^{2}$ yields $\frac{\mu}{2}\left\|\bar{y}-x_{p}^{*}\right\|^{2}\leq\left\|x_{p}^{*}-\bar{y}\right\|\nu,$ which implies that
$\left\|x_{p}^{*}-\bar{y}\right\|\leq\frac{2}{\mu}\nu\ .$ (4.4)
Putting (4.3) and (4.4) together, we have $f\left(\bar{y}\right)-f^{*}\leq{2}\nu^{2}/{\mu}$.
Summarizing two cases completes our proof. ∎
Lemma 4.2 establishes that the W-certificate provides a rigorous and practically implementable criterion for certifying the near-optimality of any candidate point $\bar{y}\in X$. Building on the bundle-based iterative framework for constructing a W-certificate for PWS functions developed in [74], we introduce Algorithm 3 below. This algorithm incorporates an accelerated prox-level approach, substantially reducing the number of oracle calls needed to generate a certificate with quality that matches or even exceeds that of the conventional BL-based method.
Algorithm 3 requires the following inputs: the convex objective $f$, a candidate solution $\bar{y}$, the estimated optimality gap $\Delta$, the desired number of supporting cuts $m$, a maximal radius $\iota_{\max}$, and a scaling parameter $\beta>0$. At each inner iteration of the One-Step subroutine (Algorithm 1, Line 6 of Algorithm 3), the algorithm projects the reference point $\bar{y}$ onto the bundle level set to generate a new point. Unlike the non-accelerated variant in [74], the evaluation points for W-certificate is selected from sequence $\left\{\underline{x}^{t,i}\right\}$ in One-Step but not $\left\{x^{t,i}\right\}$, which is because the bundle cuts are constructed by points $\left\{\underline{x}^{t,i}\right\}$. Moreover, it is worth noting that by choosing the level set parameter $\tilde{l}=f\left(\bar{y}\right)-\left(1+\beta\right)\Delta$ below the unknown optimal value $f^{*}$, we encourage exploration of different smooth pieces. As the algorithm progresses, the distance $\left\|x^{t+1,0}-\bar{y}\right\|$ from the original point $\bar{y}$ increases monotonically, as proved formally in Proposition 3.1. As soon as this distance exceeds the prescribed radius $\iota_{\max}$, or if the projection subproblem becomes infeasible (equivalently, if the distance tends to $\infty$), the maximum-of-linear-supports model $\psi\left(x\right)$ built on the evaluation set $\mathcal{P}$ is uniformly bounded below by $\tilde{l}$ within the ball $\mathcal{B}\left(\bar{y},\iota_{\max}\right)$. At this point (Line 6 of Algorithm 3), the algorithm terminates and returns $\mathcal{P}$ as an $\left(\iota,\nu\right)$-W-certificate. If the input gap estimate $\Delta$ satisfies $f\left(\bar{y}\right)-f^{*}\leq\Delta$, then Algorithm 3 is guaranteed to produce a certificate for $\bar{y}$. However, if the procedure instead exits at Line 10, which indicates the pre-assumption that $f\left(\bar{y}\right)-f^{*}\leq\Delta$ is wrong. This result is stated in Proposition 4.1 formally.
Algorithm 3 Accelerated W-certificate Generation, $\mathcal{AWG}\left(f,\bar{y},\Delta,m,\iota_{\max},\beta\right)$
1:Initialize: $x^{0}=\bar{y},\hat{x}^{0}=\bar{y},\tilde{l}=f\left(\bar{y}\right)-\left(1+\beta\right)\Delta$, $\mathcal{P}=\left\{\bar{y}\right\}$
2:for $t=1,\ldots$ do
3:  $\left(\hat{x}^{t},x^{t},\text{{Flag}},\mathcal{P}_{t},L_{t},\left\|x^{t,l_{t}}-x^{t,r_{t}}\right\|\right)\leftarrow\text{One-Step}\left(f,\bar{y},\hat{x}^{t-1},x^{t-1},\tilde{l},m\right)$
4:  $\mathcal{P}=\mathcal{P}\cup\mathcal{P}_{t}$
5:  if Flag is True or $\left\|x^{t}-\bar{y}\right\|>\iota_{\max}$ then
6:    Return True, $\hat{x}^{t}$ and $\mathcal{P}=\mathcal{P}\cap{\cal B}\left(\bar{y};\iota_{\max}\right)$ as the $\left(\iota_{\max},\left(1+\beta\right)\Delta/\iota_{\max}\right)$-W-certificate for point $\bar{y}$
7:  end if
8:  Compute $\bar{L}\left(t\right)$ (defined in (3.2)) based on $\left\|x^{t,l_{t}}-x^{t,r_{t}}\right\|$
9:  if $\left(\bar{L}\left(t\right)\left\|x^{t}-\bar{y}\right\|^{2}<\omega_{t}\beta\Delta-3\left(1+\beta\right)\Delta\right)$ or $t\geq\sqrt{\left(2\iota_{\max}^{2}\bar{L}\left(t\right)+6\left(1+\beta\right)\Delta\right)/\beta\Delta}$ then
10:    Return $\text{{\text{False}}},\hat{x}^{t},\ldots$
11:  end if
12:end for
In order to formalize the fundamental property of Algorithm 3, we begin by isolating and stating Lemma 4.3, which will serve as a recurring tool in the subsequent analysis. Building on this lemma, we then establish Proposition 4.1, which offers a rigorous characterization of the behavior of Algorithm 3.
Lemma 4.3.
Suppose Algorithm 3 has generated the point set $\mathcal{P}_{t}$ and the iterate $x^{t}$ at outer iteration $t$. For any $\iota\in(0,\|x^{t}-\bar{y}\|)$, define $\tilde{\mathcal{P}}(\iota):=\mathcal{P}_{t}\cap\mathcal{B}(\bar{y},\iota)$. Then we have $\min_{x\in\mathcal{B}(\bar{y},\iota)}\psi_{\tilde{\mathcal{P}}(\iota),f}(x)>\tilde{l}$.
Proof.
Denote the set that generated by all pass linear cut as
$\Psi_{t,f}=\left\{x:\psi_{\mathcal{P}_{t},f}\left(\hat{x}\right)\leq\tilde{l}\right\},\text{ where }\mathcal{P}_{t}=\left\{\underline{x}^{1,1},\ldots,\underline{x}^{1,m},\ldots,\underline{x}^{t-1,1},\ldots,\underline{x}^{t-1,m}\right\}\ ,$ (4.5)
then for every $\hat{x}\in\Psi_{t,f}\cap X$, we claim $\left\langle\hat{x}-x^{t+1,0},x^{t+1,0}-\bar{y}\right\rangle\geq 0$ holds for any $t\geq 1$.
We prove this by induction. For base case $t=1$, we have $x^{2,0}=x^{1,m}$. Since $x^{1,0}=\bar{y}$, then $\bar{X}\left(t,0\right)$ in option I is eliminated. Hence, $x^{1,m}$ is the closest point to set $\left\{x:\psi_{\mathcal{P}_{1},f}\left(x\right)\leq\tilde{l}\right\}$, which implies that for any point satisfies $\hat{x}\in\Psi_{1,f}\cap X$, we have $\left\langle\hat{x}-x^{t+1,0},x^{t+1,0}-\bar{y}\right\rangle\geq 0$. Suppose the claim (4.5) holds for $1,\ldots,t$, we aim to prove if every $\hat{x}\in\Psi_{t,f}\cap X$, we have $\left\langle\hat{x}-x^{t+1,0},x^{t+1,0}-\bar{y}\right\rangle\geq 0$. Since $\hat{x}\in\Psi_{t-1,f}$ implies $\left\langle\hat{x}-x^{t,0},x^{t,0}-\bar{y}\right\rangle\geq 0$ by induction. Hence, the goal of our proof is equivalent to for any $\hat{x}\in X$, $\left\langle\hat{x}-x^{t,0},x^{t,0}-\bar{y}\right\rangle\geq 0$ and $\psi_{\mathcal{P}_{t}\backslash\mathcal{P}_{t-1},f}\left(\hat{x}\right)\leq\tilde{l}$, we have $\left\langle\hat{x}-x^{t+1,0},x^{t+1,0}-\bar{y}\right\rangle\geq 0$. Note that the region
$\left\{x:\left\langle\hat{x}-x^{t,0},x^{t,0}-\bar{y}\right\rangle\geq 0,\psi_{\mathcal{P}_{t}\backslash\mathcal{P}_{t-1},f}\left(\hat{x}\right)\leq\tilde{l}\right\}$
of $\hat{x}$ is exactly the feasible region $X(t,m)$ of last bundle subproblem at $t$-th iteration. Since $x^{t+1,0}=x^{t,m}$ is the closest point to the set $X(t,m)$, then we have any $\hat{x}\in X(t,m)$, $\left\langle\hat{x}-x^{t+1,0},x^{t+1,0}-\bar{y}\right\rangle\geq 0$ holds.
Now, we are ready to prove the main result of Lemma 4.3. Since sequence $\left\{\left\|x^{t,i}-\bar{y}\right\|\right\}$ is non-decreasing by Proposition 3.1 and $\left\|x^{1,0}-\bar{y}\right\|=0$. Note that $x^{i,m}=x^{i+1,0}$ holds for any $i\in\left[t-1\right]$. Hence there exists a index pair $\left(\tilde{t},j\right)\in\left[t\right]\times\left[m\right]$ such that $\iota\in\left(\left\|x^{\tilde{t},j-1}-\bar{y}\right\|,\left\|x^{\tilde{t},j}-\bar{y}\right\|\right]$. We first denote $\mathcal{P}\left(\tilde{t},j\right)=\left\{\underline{x}^{\tilde{t},1},\ldots,\underline{x}^{\tilde{t},j}\right\}$, then the set $X\left(t,i\right)$ in Algorithm 1 can be rewritten as $X\left(\tilde{t},j\right)=\left\{x:\psi_{\mathcal{P}\left(\tilde{t},j\right),f}\left(x\right)\leq\tilde{l}\ ,\ \left\langle x-x^{\tilde{t},0},x^{\tilde{t},0}-\bar{y}\right\rangle\geq 0\right\}$. Since $x^{\tilde{t},j}$ is the closest point to set $X\left(\tilde{t},j\right)$. Hence, $\forall\hat{x}\in X\left(\tilde{t},j\right)\cap X$, we have $\left\|\hat{x}-\bar{y}\right\|\geq\left\|x^{\tilde{t},j}-\bar{y}\right\|\geq\iota$. This implies that for any $\hat{x}\in\mathcal{B}\left(\bar{y};\iota\right)\cap X$, we have $\psi_{\mathcal{P}\left(\tilde{t},j\right),f}\left(\hat{x}\right)>\tilde{l}$ or $\left\langle\hat{x}-x^{\tilde{t},0},x^{\tilde{t},0}-\bar{y}\right\rangle<0$. It follows from every $\hat{x}\in\Psi_{t,f}\cap X$ implies $\left\langle\hat{x}-x^{t+1,0},x^{t+1,0}-\bar{y}\right\rangle\geq 0$ that $\left\langle\hat{x}-x^{\tilde{t},0},x^{\tilde{t},0}-\bar{y}\right\rangle<0$ implies $\psi_{\mathcal{P}_{\tilde{t}},f}\left(\hat{x}\right)>\tilde{l}$. Hence, we complete the proof. ∎
Proposition 4.1.
Let the weighted sequence $\omega_{t}=\left(t+2\right)\left(t+3\right)/2$ and $\alpha_{t}=4/\left(t+3\right)$ in Algorithm 3. If the objective function $f$ is convex and $f\left(\bar{y}\right)-f^{*}\leq\Delta$, the Algorithm 3 will return $\left(\iota_{\max},\left(1+\beta\right)\Delta/\iota_{\max}\right)$-W-certificate at most $\left(m\sqrt{\left(2\iota_{\max}\bar{L}\left(t\right)+6\left(1+\beta\right)\Delta\right)/\left(\beta\Delta\right)}\right)$ gradient calls in total.
Proof.
Algorithm 3 is a special case of APEX, with $\tilde{l}=f(\bar{y})-(1+\beta)\Delta$, $\alpha_{t}=4/(t+3)$, and $\omega_{t}=(t+2)(t+3)/2$. We now show that when Algorithm 3 reaches Line 6, the set $\mathcal{P}$ forms a valid $\left(\iota_{\max},{(1+\beta)\Delta}/{\iota_{\max}}\right)$-W-certificate for $\bar{y}$. By construction, $\mathcal{P}\subseteq\mathcal{B}(\bar{y};\iota_{\max})$. It remains to bound the normalized Wolfe gap $\mathcal{V}_{\mathcal{P},f}(\iota_{\max};\bar{y})$:
$\begin{aligned} &\mathcal{V}_{\mathcal{P},f}\left(\iota_{\max};\bar{y}\right)=\frac{1}{\iota_{\max}}\max_{x\in{\cal B}\left(\bar{y};\iota_{\max}\right)}\left\{\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(x\right)\right\}\\ &\overset{(a)}{\leq}\frac{1}{\iota_{\max}}\max_{x\in{\cal B}\left(\bar{y};\iota_{\max}\right)}\left\{f\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(x\right)\right\}\overset{(b)}{<}\frac{1}{\iota_{\max}}\left\{f\left(\bar{y}\right)-\tilde{l}\right\}=\frac{\left(1+\beta\right)\Delta}{\iota_{\max}}\end{aligned}\ ,$
where $(a)$ holds by $f$ is convex and $(b)$ holds by Lemma 4.3. In the subsequent, we show that if $f\left(\bar{y}\right)-f^{*}\leq\Delta$ holds, the Algorithm 3 will enter Line 6 and the total gradient calls would not exceed $\left(m\sqrt{\left(2\iota_{\max}\bar{L}\left(t\right)+6\left(1+\beta\right)\Delta\right)/\beta\Delta}\right)$. It follows from Proposition 3.1 that
$\left(t+2\right)\left(t+3\right)\left(f\left(\hat{x}^{t}\right)-\left(f\left(\bar{y}\right)-\left(1+\beta\right)\Delta\right)\right)/2-3\left(1+\beta\right)\Delta\leq\bar{L}\left(t\right)\left\|x^{t,0}-\bar{y}\right\|^{2}\ .$
Combining the above result and $f\left(\bar{y}\right)-f^{*}\leq\Delta$ gives
$\left(t+2\right)\left(t+3\right)\beta\Delta/2-3\left(1+\beta\right)\Delta\leq\bar{L}\left(t\right)\left\|x^{t,0}-\bar{y}\right\|^{2}\ .$
If $t\geq\sqrt{\left(2\iota_{\max}^{2}\bar{L}\left(t\right)+6\left(1+\beta\right)\Delta\right)/\beta\Delta}$, then we must have $\left\|x^{t,0}-\bar{y}\right\|\geq\iota_{\max}$, which implies this result. ∎
We make a few remarks regarding the result. First, although algorithms such as the bundle-level method proposed in [74] or $\mathcal{AWG}$ are useful for constructing certificates, it is important to emphasize that W-certificate itself is independent of any specific algorithm. In other words, W-certificate is an inherent property of a candidate point, and an algorithm merely serves as a mechanism to generate a point set that qualifies as a W-certificate for that candidate. Second, W-certificate is also independent of the $\mu$-QG condition. If the goal is to construct an upper bound on the optimality gap, then a QG assumption is indeed required (see Lemma 4.2). However, if W-certificate is adopted purely as a termination criterion (like gradient norm), the QG condition is not necessary.
5 Restarted APEX for known $\mu^{*}$
By leveraging the W-certificate, we can now address convex PWS optimization under a $\mu^{*}$-QG without requiring prior knowledge of the optimal value $f^{*}$. We maintain two evolving estimates: one is a lower bound ($\bar{f}-\Delta$) on $f^{*}$ obtained from W-certificate and the other is an upper bound ($\bar{f}$) on $f^{*}$ tracked by the best objective value encountered along the iterates. We then restrict our search to the level between these two bounds. Algorithm 4 implements this idea via a simple restart mechanism. Algorithm 4 first initializes a trustable optimality gap $\Delta_{0}$ by $\mu$-QG roughly. Subsequently, Algorithm 4 sets the level between the current upper bound and lower bound and run the accelerated bundle level for piecewise smooth until one of two restarted conditions is met:
  • Upper bound improvement (Line 8): the upper bound has a good improvement, specifically, the gap between upper bound and previous lower bound decrease by a factor $\theta$.
  • Lower bound improvement (Line 10): an iterate $x_{s}^{t}$ attains $\left\|x_{s}^{t}-\bar{y}\right\|$ large enough to generate a new W-certificate, which by Lemma 4.2 yields a strictly tighter lower bound $(\bar{f}-\theta\Delta)$ on $f^{*}$.
By alternately tightening the upper and lower bounds through these two triggers, Algorithm 4 converges to $f^{*}$ without requiring explicit knowledge of $f^{*}$.
Algorithm 4 Restarted APEX (rAPEX) for known $\mu^{*}$
1:objective function $f$, the initial point $\bar{y}^{0}$, QG modulus $\mu^{*}$, the number of cuts $m$, lower bound $\underline{f}^{0}$ of optimal value $f^{*}$, shrink factor $\theta\in\left(\frac{1}{2},1\right)$
2:Initialize: upper bound $\bar{f}^{0}=f\left(\bar{y}^{0}\right)$ of optimal value $f^{*}$, optimality gap $\Delta_{0}=\frac{\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}}{2\mu^{*}}$, lower bound $\underline{f}^{0}=\bar{f}^{0}-\Delta_{0}$, $s=0$
3:while True do
4:   $s=s+1,x_{s}^{1,0}=\bar{y}^{s-1},\hat{x}_{s}^{0}=\bar{y}^{s-1},\tilde{l}^{s-1}=\bar{f}^{s-1}-\theta\Delta_{s-1}$
5:  for $t=1,\ldots$ do
6:   $\left(\hat{x}_{s}^{t},x_{s}^{t},\text{$\text{{Flag}}$},\mathcal{P}_{s},\ldots\right)\leftarrow\text{One-Step}\left(f,\bar{y}^{s-1},\hat{x}_{s}^{t-1},x_{s}^{t-1},\tilde{l}^{s-1},m\right)$
7:   if $f\left(\hat{x}^{t}\right)-\underline{f}^{s-1}\leq\theta\Delta_{s-1}$ then
8:      $\bar{y}^{s}=\hat{x}^{t},\bar{f}^{s}=f\left(\hat{x}^{t}\right),\Delta_{s}=\theta\Delta_{s-1},\underline{f}^{s}=\underline{f}^{s-1}$ and go to Line 4
9:   else if $\left\|x_{s}^{t}-\bar{y}\right\|^{2}>\frac{2\theta\Delta_{s-1}}{\mu}$ or Flag is True then
10:      $\bar{y}^{s}=\hat{x}^{t},\bar{f}^{s}=f\left(\hat{x}^{t}\right),\Delta_{s}=\theta\Delta_{s-1},\underline{f}^{s}=\bar{f}^{s-1}-\Delta_{s}$ and go to Line 4
11:   end if
12:  end for
13:end while
In the following Theorem 5.1, we state the convergence result of Algorithm 4 formally.
Theorem 5.1.
Suppose function $f$ is a convex function which satisfies the quadratic growth condition with modulus $\mu$, and let $\omega_{t}=\left(t+2\right)\left(t+3\right)/2$ and $\alpha_{t}=4/\left(t+3\right)$ then the following statements for Algorithm 4 hold:
  1. 1.
    If Algorithm 4 reaches Line 10, then $\mathcal{P}_{s}$ is a $\left(\sqrt{2\theta\Delta_{s-1}/\mu^{*}},\sqrt{\mu^{*}\theta\Delta_{s-1}/2}\right)$-W-certificate for point $\bar{y}^{s-1}$ in function $f$.
  2. 2.
    If $L_{\text{avg}}=\max_{s=1,\ldots}L\left(\bar{t}_{s}\right)$ (defined in (3.2)) of approximate Lipschitz constant $L_{t}$ (defined in (3.2)) across every outer loop, where $\bar{t}_{s}$ is the number of inner loop iterations consumed in each $s$-th outer loop, then to obtain an $\varepsilon$-optimal solution, $\left(f\left(x\right)-f^{*}\leq\varepsilon\right)$, the overall gradient evaluation calls is at most $O(1)\left(m\cdot\sqrt{{L_{\textrm{avg}}}/{\mu^{*}}}\log\left(\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}/{2\mu^{*}\varepsilon}\right)\right)$.
  3. 3.
    If $f$ is also a $\left(k,L\right)$-PWS function with $k\leq m$, then $L_{\text{avg}}=O(1)L$ in the above complexity, i.e., to obtain an $\varepsilon$-optimal solution $\left(f\left(x\right)-f^{*}\leq\varepsilon\right)$, the overall gradient evaluation calls is at most $O(1)\left(m\cdot\sqrt{{L}/{\mu^{*}}}\log\left({\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}}/{2\mu^{*}\varepsilon}\right)\right)$.
Proof.
Part 1. Note Algorithm 4 is similar to Algorithm 3. The two Algorithms run the same One-Step subroutine. Hence, By the similar proof procedure of Proposition 4.1 with $(1+\beta)\Delta=\theta\Delta_{s-1},\iota_{\max}=\sqrt{2\theta\Delta_{s-1}/\mu^{*}}$ that Algorithm 4 reaching Line 10 will generate point set $\mathcal{P}_{s}\cap\mathcal{B}\left(\bar{y}^{s-1};\sqrt{\theta\Delta_{s-1}/\left(2\mu^{*}\right)}\right)$ as $\left(\sqrt{2\theta\Delta_{s-1}/\mu^{*}},\sqrt{\mu^{*}\theta\Delta_{s-1}/2}\right)$-W-certificate for point $\bar{y}^{s-1}$ in function $f$.
Part 2. We divide the argument into two principal parts: 1. Validity of the lower bound: every updating of lower bound $\underline{f}^{s}$ is valid, i.e., $\underline{f}^{s}\leq f^{*}.$ 2. Per-restart oracle cost: each individual restart invokes the accelerated bundle level subroutine only a constant number of times, and thus incur the final complexity result since the gap shrinks by at least the constant factor $\theta$.
We first prove that Line 10 generates a proper lower bound satisfying $\underline{f}^{s}\leq f^{*}.$ By Lemma 4.2, $\left(\sqrt{2\theta\Delta_{s-1}/\mu^{*}},\sqrt{\mu^{*}\theta\Delta_{s-1}/2}\right)$-W-certificate implies a valid gap: $f\left(\bar{y}^{s-1}\right)-f^{*}\leq\theta\Delta_{s-1}$. Hence, we can set $\underline{f}^{s}=\bar{f}^{s-1}-\theta\Delta_{s-1}\leq f^{*}$ in Line 10 of Algorithm 4.
In the following, we proceed by contradiction to prove that at most
$\bar{t}=\left\lceil\sqrt{\frac{6\theta+{4L_{\text{avg}}\theta}/{\mu}}{\left(2\theta-1\right)}}\right\rceil$ (5.1)
iterations, the Algorithm 4 will enter Line 8 or Line 10 to restart algorithm. Suppose on the contrary $f\left(\hat{x}_{s}^{\bar{t}}\right)-\underline{f}^{s-1}>\theta\Delta_{s-1}$. Due to the non-increasing of $f\left(\hat{x}_{s}^{t}\right)$ in One-Step (Algorithm 1), we have $f\left(\hat{x}_{s}^{t}\right)-\underline{f}^{s-1}>\theta\Delta_{s-1},\forall t\in\left[\bar{t}\right]$, which implies that $f\left(\hat{x}_{s}^{t}\right)-\tilde{l}^{s-1}\geq\left(2\theta-1\right)\Delta_{s-1},\forall t\in\left[\bar{t}\right]$. If there exists an iteration smaller than $\bar{t}$ that the subproblem is infeasible, then our proof has finished. Hence, without loss of generality, we assume all subproblems are feasible. Then it follows from Proposition 3.1 that
$\begin{split}\omega_{\bar{t}}\left(\left(2\theta-1\right)\Delta_{s-1}\right)-\omega_{1}\left(1-\frac{\alpha_{1}}{2}\right)\theta\Delta_{s-1}&\overset{(a)}{<}\omega_{\bar{t}}\left(f\left(\hat{x}_{s}^{t}\right)-\tilde{l}^{s-1}\right)-\omega_{1}\left(1-\frac{\alpha_{1}}{2}\right)\left(f\left(\hat{x}_{s}^{1}\right)-\tilde{l}^{s-1}\right)\\ &\leq\bar{L}\left(\bar{t}\right)\left\|x^{N}-\bar{y}\right\|^{2}\ ,\end{split}$ (5.2)
where $(a)$ holds by $f\left(\hat{x}_{s}^{t}\right)>\underline{f}^{s-1}+\theta\Delta_{s-1}$ and $\tilde{l}^{s-1}=\bar{f}^{s-1}-\theta\Delta_{s-1}$. If follows from the definition of $\tilde{l}^{s-1}$ and non-increasing of $f\left(\hat{x}_{s}^{t}\right)$ that
$f\left(\hat{x}_{s}^{t-1}\right)-\tilde{l}^{s-1}\leq f\left(\bar{y}^{s-1}\right)-\tilde{l}^{s-1}=\theta\Delta_{s-1}\ .$
Hence, combining $\omega_{1}\left(1-\alpha_{1}/2\right)=3$, (5.2), $\bar{L}\left(\bar{t}\right)\leq L_{\text{avg}}$ and (3.12) gives
$\begin{split}\left\|x^{\bar{t}}-\bar{y}\right\|^{2}&\geq\frac{\omega_{\bar{t}}\left(\left(2\theta-1\right)\Delta_{s-1}\right)-\omega_{1}\left(1-\frac{\alpha_{1}}{2}\right)\theta\Delta_{s-1}}{L_{\text{avg}}}\geq\frac{\left(\omega_{\bar{t}}\left(2\theta-1\right)-6\theta\right)\Delta_{s-1}}{L_{\text{avg}}}\\ &=\frac{\left(\frac{\left(\bar{t}+2\right)\left(\bar{t}+3\right)}{2}\left(2\theta-1\right)-3\theta\right)\Delta_{s-1}}{L_{\text{avg}}}\geq\frac{\left(\frac{1}{2}\left(2\theta-1\right)\bar{t}^{2}-3\theta\right)\Delta_{s-1}}{L_{\text{avg}}}\overset{(a)}{\geq}\frac{2\theta\Delta_{s-1}}{\mu^{*}}\end{split}\ ,$ (5.3)
where $(a)$ holds by the definition of $\bar{t}$. Hence, we complete the proof that there exists a time stamp $t\in\left[\bar{t}\right]$ such that Algorithm 4 will enter Line 8 or 10. No matter which of the three conditions in Line 8 Algorithm 4 achieves, it follows from the non-increasing property of $f\left(\hat{x}_{s}^{t}\right)$ that $\Delta_{s}\leq\theta\Delta_{s-1}$. Hence, we have $f\left(\bar{y}^{s}\right)-f^{*}\leq\theta^{s}\Delta_{0}=\theta^{s}\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}/\left(2\mu^{*}\right)$ and for an $\varepsilon$-optimal solution, the gradient evaluation is at most $O(1)\left(m\cdot\sqrt{\frac{L_{\textrm{avg}}}{\mu^{*}}}\log\frac{\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}}{2\mu^{*}\varepsilon}\right)$ in total.
Part 3. At each $s$-th phase, the objective would not change very large. Hence, the term $\bar{L}\left(\bar{t}\right)$ can be bounded as follows:
$\begin{split}\bar{L}\left(\bar{t}\right)&\leq\frac{3\sum_{t=1}^{\bar{t}}\mathbf{1}_{\mathcal{E}(t)}\omega_{t}\left(\alpha_{t}\left(f\left(\hat{x}_{s}^{t-1}\right)-\tilde{l}^{s-1}\right)\right)}{\sum_{t=1}^{\bar{t}}\mathbf{1}_{\mathcal{E}(t)}\left(f\left(\hat{x}_{s}^{t-1}\right)-\tilde{l}^{s-1}\right)\Big/\left(\alpha_{t}\cdot L_{t}\right)}\overset{(a)}{\leq}\frac{3\theta\Delta_{s-1}\sum_{t=1}^{\bar{t}}\mathbf{1}_{\mathcal{E}(t)}\omega_{t}\alpha_{t}}{\left(2\theta-1\right)\Delta_{s-1}\sum_{t=1}^{\bar{t}}\mathbf{1}_{\mathcal{E}(t)}\Big/\left(\alpha_{t}\cdot L_{t}\right)}\ ,\\ &\overset{(b)}{\leq}\frac{24\theta\sum_{t=1}^{\bar{t}}\mathbf{1}_{\mathcal{E}(t)}\alpha_{t}^{-1}}{\left(2\theta-1\right)\sum_{t=1}^{\bar{t}}\mathbf{1}_{\mathcal{E}(t)}\alpha_{t}^{-1}L_{t}^{-1}}\end{split}$ (5.4)
where $(a)$ follows from $\theta\Delta_{s-1}\geq f\left(\hat{x}_{s}^{t}\right)-\tilde{l}^{s-1}\geq\left(2\theta-1\right)\Delta_{s-1}$ holds by $\omega_{t}\alpha_{t}\leq 8\alpha_{t}^{-1}$. If function $f$ is $\left(k,L\right)$-PWS function with $k\leq m$, then we have $L_{t}\leq L$. Hence, $\bar{L}\left(\bar{t}\right)=O(1)L$. Combining this result and complexity in Part 2 yields the result in Part 3. ∎
A couple of remarks are in order regarding the convergence results above. First, for a prescribed accuracy $\varepsilon$, the number of oracle evaluations required to obtain an $\varepsilon$ optimal solution is $O(1)\left(m\cdot\sqrt{L_{\textrm{avg}}/\mu^{*}}\log\left({1}/{\varepsilon}\right)\right)$. Once the bundle of supporting cuts fully captures the underlying piecewise smooth structure so that $L_{\text{avg}}=O(1)L$, with $L$ being the true global Lipschitz constant (see Proposition 3.2 and the corresponding remark for details), this rate matches, up to the multiplicative factor $m$, the optimal complexity bound for smooth convex optimization under the $\mu^{*}$-QG assumption. Moreover, since the number of active pieces may evolve as the algorithm progresses, the method naturally transitions between a coarse piecewise smooth regime and a smooth regime, adaptively accelerating convergence once the structural decomposition is well resolved. Second, Algorithm 4 requires only knowledge of the QG constant $\mu^{*}$. It does not depend on the global Lipschitz constant $L$, nor does it require prior knowledge of the exact number of active pieces $k$.
6 Restarted APEX for unknown $\mu^{*}$
In this section, we propose an algorithm that eliminates the requirement for a priori knowledge of the QG constant $\mu^{*}$. To achieve this, we embed Algorithm 4 within a systematic ”guess–and–check” mechanism, yielding Algorithm 6. At each outer iteration, we maintain an estimate $\hat{\mu}_{s}>0$ of the true QG constant $\mu^{*}$ and use it to define a candidate optimality gap $\tilde{\Delta}_{s}$, which governs two verification mechanisms: $(1)$ Within a fixed budget of inner iterations, $\mathcal{AWG}$ attempts to produce a W-certificate for the current reference point $\bar{y}$. Failure to generate such a certificate implies that $\tilde{\Delta}_{s}$ underestimates the true optimality gap $f\left(\bar{y}^{s-1}\right)-f^{*}$, which, under the QG assumption, can only occur if $\hat{\mu}_{s}>\mu^{*}$. In that case (Line 6 of Algorithm 4), $\hat{\mu}_{s}$ is halved and the guess–and–check loop is restarted. $(2)$ If $\mathcal{AWG}$ succeeds, we invoke $\mathcal{AGR}$ to simultaneously refine the upper bound of $f\left(x\right)$ and the lower bound estimate. Similar to Algorithm 4, $\mathcal{AGR}$ will make progress (upper bound or lower bound) before exhausting its budget if $f\left(\bar{y}^{s-1}\right)-f^{*}\leq\tilde{\Delta}_{s}$ holds (see Proposition 6.1 for more details).
Algorithm 5 Accelerated Gap Reduction, $\mathcal{AGR}\left(f,\underline{f},\Delta,\bar{y},\mu,\theta,m\right)$
1:Initialize: $\tilde{l}=f\left(\bar{y}\right)-\theta\Delta,x^{0}=\bar{y}$
2:for $t=1,\ldots$ do
[…]
4:  Compute $\bar{L}\left(t\right)$ (defined in (3.2)) based on $\left\|x^{t,l_{t}}-x^{t,r_{t}}\right\|$ and $\mathcal{P}=\mathcal{P}\cup\mathcal{P}_{t}$
5:  if Flag is True or $\left\|x^{t}-\bar{y}\right\|^{2}\geq\frac{2\theta\Delta}{\mu}$ then
6:    Return $\hat{x}^{t},\text{{True}},\mathcal{P}\cap\mathcal{B}\left(\bar{y},\sqrt{\frac{2\theta\Delta}{\mu}}\right)$
7:  else if $f\left(\hat{x}^{t}\right)-\underline{f}\leq\theta\Delta$ then
8:    Return $\hat{x}^{t},\text{{\text{False}}},\ldots$
9:  end if
10:end for
Algorithm 6 rAPEX for $\mu$ is unknown
1:objective function $f$, the initial point $\bar{y}^{0}$, strong convexity estimation $\hat{\mu}_{1}$, the number of cuts $m$, shrink factor $\theta\in\left(\frac{1}{2},1\right)$, extra factor $\beta>0$
2:Initialize: upper bound $\bar{f}=f\left(\bar{y}^{0}\right)$, $\tilde{\Delta}_{1}=\frac{2\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}}{\hat{\mu}_{1}}\ ,\ \underline{f}=\bar{f}-\tilde{\Delta}_{1}\ ,\ \mu_{0}=1\ ,\ {\Delta}_{0}=+\infty$
3:for $s=1,\ldots$ do
4:   $\left(\text{{Quar-Flag}},\hat{x}^{s},\tilde{\mathcal{P}}_{s},\ldots\right)\leftarrow\mathcal{AWG}\left(f\left(\cdot\right),\bar{y}^{s-1},\tilde{\Delta}_{s},m,\sqrt{\frac{2\left(1+\beta\right)\tilde{\Delta}_{s}}{\hat{\mu}_{s}}},\beta\right)$
5:  if Quar-Flag is False then
6:    $\hat{\mu}_{s}=\hat{\mu}_{s}/4\ ,\ \tilde{\Delta}_{s}=\min\left\{\frac{9\mu_{s-1}}{4\hat{\mu}_{s}}\Delta_{s-1},\frac{2}{\hat{\mu}_{s}}\left\|f^{\prime}\left(\bar{y}^{s-1}\right)\right\|^{2}\right\}\ ,\ \underline{f}=\bar{f}-\tilde{\Delta}_{s}$ and go to Line 4
7:  end if
8:   $\Delta_{s}=\left(1+\beta\right)\tilde{\Delta}_{s}\ ,\ \mu_{s}=\hat{\mu}_{s}\ ,\ \iota_{s}=\sqrt{\frac{2\Delta_{s}}{\mu_{s}}}\ ,\ \mathcal{P}_{s}=\tilde{\mathcal{P}}_{s}$
9:   $\left(\hat{x}^{s},\text{{Lower-Flag}},\tilde{\mathcal{P}}_{s}\right)\leftarrow\mathcal{AGR}\left(f\left(\cdot\right),\underline{f},\tilde{\Delta}_{s},\bar{y}^{s-1},\hat{\mu}_{s},\theta,m\right)$
10:  if Lower-Flag is True then
11:    $\tilde{\Delta}_{s}=\theta\tilde{\Delta}_{s},\underline{f}=\bar{f}-\tilde{\Delta}_{s},\Delta_{s}=\tilde{\Delta}_{s},\mu_{s}=\hat{\mu}_{s},\iota_{s}=\sqrt{\frac{2\Delta_{s}}{\mu_{s}}},\mathcal{P}_{s}=\tilde{\mathcal{P}}_{s}$ and go to Line 9
12:  end if
13:  if $f(\hat{x}^{s})<\underline{f}$ then
14:    go to Line 6
15:  end if
16:   $\bar{y}^{s}=\hat{x}^{s},\bar{f}=f\left(\bar{y}^{s}\right),\tilde{\Delta}_{s+1}=\bar{f}-\underline{f}$
17:end for
Beyond the “guess-and-check” procedure for validating $\hat{\mu}_{s}$, Algorithm 6 crucially relies on a certificate-transfer mechanism to carry forward valid W-certificate and thus maintain sharp lower bounds on $f^{*}$ even the QG modulus is misspecified. Concretely, suppose at the $s$-th outer stage $(s>1)$, Algorithm 6 enters Line 6, so that $\mathcal{AWG}$ fails to produce a new W-certificate for the current iterate $\bar{y}^{s-1}$. Although no fresh W-certificate for $\bar{y}^{s-1}$, we still hold that the most recent valid $\left(\iota_{s-1},\nu_{s-1}\right)$-W-certificate that was generated for the previous iterate $\bar{y}^{s-2}$. In the following Lemma 6.1, we show that a W-certificate of $\bar{y}^{s-1}$ can be transferred from $\left(\iota_{s-1},\nu_{s-1}\right)$-W-certificate of $\bar{y}^{s-2}$. By invoking this transferred certificate, we obtain a valid and typically much tighter–lower bound on $f^{*}$ than the coarse estimate $f\left(\bar{y}^{s-1}\right)-\left\|f^{\prime}\left(\bar{y}^{s-1}\right)\right\|^{2}\big/\left(2\hat{\mu}_{s}\right)$, thereby preserving the accelerated convergence rate.
This strategy is for ensuring that all lower-bound estimates remain as tight as possible throughout the algorithm.
We first show that an $\left(\iota,\nu\right)$W-certificate established for one candidate solution can be transferred to another point under suitable conditions in the following Lemma 6.1.
Lemma 6.1.
[…]
Proof.
[…]
$\begin{split}\tilde{\iota}\mathcal{V}_{\tilde{\mathcal{P}},f}\left(\tilde{\iota};\tilde{y}\right)&=\max_{x\in\mathcal{B}\left(\tilde{y};\tilde{\iota}\right)}\left\{\psi_{\mathcal{P},f}\left(\tilde{y}\right)-\psi_{\mathcal{P},f}\left(x\right)\right\}\\ &\overset{(a)}{\leq}\psi_{\mathcal{P},f}\left(\tilde{y}\right)-\psi_{\mathcal{P},f}\left(\bar{y}\right)+\max_{x\in\mathcal{B}\left(\bar{y};\left(1+2c\right)\iota\right)}\left\{\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(x\right)\right\}\\ &=\psi_{\mathcal{P},f}\left(\tilde{y}\right)-\psi_{\mathcal{P},f}\left(\bar{y}\right)+\left(1+2c\right)\iota\mathcal{V}_{\mathcal{P},f}\left(\left(1+2c\right)\iota;\bar{y}\right)\\ &\overset{(b)}{\leq}\psi_{\mathcal{P},f}\left(\tilde{y}\right)-\psi_{\mathcal{P},f}\left(\bar{y}\right)+\left(1+2c\right)\iota\mathcal{V}_{\mathcal{P},f}\left(\iota;\bar{y}\right)\\ &=\psi_{\mathcal{P},f}\left(\tilde{y}\right)-\psi_{\mathcal{P},f}\left(\bar{y}\right)+\left(1+2c\right)\max_{x\in\mathcal{B}\left(\tilde{y};\iota\right)}\left\{\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(x\right)\right\}\\ &\overset{(c)}{\leq}f\left(\tilde{y}\right)+cf\left(\bar{y}\right)-\left(1+2c\right)\tilde{l}\end{split}$ (6.2)
Here, $(a)$ holds by taking the definition of $\tilde{\iota}$ and $\left\|\bar{y}-\tilde{y}\right\|\leq c\iota$, $(b)$ follows from $\mathcal{V}_{\mathcal{P},f}\left(\left(1+2c\right)\iota;\bar{y}\right)\leq\mathcal{V}_{\mathcal{P},f}\left(\iota;\bar{y}\right)$ since it is a non-increasing function by Lemma 4.1 and $(c)$ holds by function $f$ is convex function and $\min_{x\in\mathcal{B}\left(\bar{y};\iota\right)\cap X}\psi_{\mathcal{P},f}\left(x\right)\geq\tilde{l}$. Hence, we complete our proof. ∎
We are now ready to present the main convergence theorem for Algorithm 6. Collectively, the preceding lemmas (Lemma 6.2 and 6.3) ensure that the lower bound estimate of $f^{*}$ remains valid under the current QG constant estimate $\hat{\mu}_{s}$. Lemma 6.4 gives the result that the upper bound of total number to quarter estimated QG $\hat{\mu}_{s}$. Proposition 6.1 gives the complexity result for subroutine $\mathcal{AGR}$. Building on this foundation, Theorem 6.1 provides the complete convergence guarantee.
Lemma 6.2.
Algorithm 6 will generate W-certificate in Line 8 or Line 11. Specifically,
  1. 1.
    If Algorithm 6 reaches Line 8, then $\mathcal{AWG}$ will generate a point set $\mathcal{P}_{s}$ as $\left(\sqrt{2\Delta_{s}/\mu_{s}},\sqrt{\mu_{s}\Delta_{s}/2}\right)$-W-certificate for point $\bar{y}^{s-1}$ in function $f\left(x\right)$.
  2. 2.
    If Algorithm 6 reaches Line 11, then $\mathcal{AGR}$ will generate a point set $\mathcal{P}_{s}$ as $\left(\sqrt{2\Delta_{s}/\mu_{s}},\sqrt{\mu_{s}\Delta_{s}/2}\right)$-W-certificate for point $\bar{y}^{s-1}$ in function $f$. Here $\Delta_{s}$ is updated newly in Line 11.
Proof.
We verify the validity of the two W-certificates in order.
Part 1: It follows that $\mathcal{AWG}$ will return a W-certificate unless it enters Line 10 in Algorithm 3, i.e., it return Quar-Flag equals to False. Hence, if Algorithm 6 runs to Line 8, the $\mathcal{AWG}$ subroutine has generated a W-certificate successfully. Specifically, by Proposition 4.1, the return point set $\mathcal{P}_{s}$ is an $\left(\iota_{\max},(1+\beta)\Delta/\iota_{\max}\right)=\left(\sqrt{2\Delta_{s}/\mu_{s}},\sqrt{\mu_{s}\Delta_{s}/2}\right)$-W-certificate for point $\bar{y}^{s-1}$.
Part 2: The proof process for Algorithm 6 enters Line 11 is similar to Proposition 4.1. ∎∎
Lemma 6.3.
Suppose function $f$ is convex satisfies $\hat{\mu}_{s}$-QG, then Algorithm 6 updates $\underline{f}$ at Lines 6 and $1$ and the lower bound $\underline{f}$ satisfying $\underline{f}\leq\min_{x\in X}f\left(x\right)$.
Proof.
The parameter $\underline{f}$ is updated in two situations (Line 6 and 11) within the algorithm. We will analyze each case to demonstrate that $\underline{f}$ is always set to a value less than or equal to the optimum $f^{*}$. The value $\tilde{\Delta}_{s}$ is chosen as the minimum of several estimates for the optimality gap $f\left(\bar{y}^{s-1}\right)-\min_{x\in X}f\left(x\right)$, so proving that all of these estimates are valid upper bounds on the gap is needed.
Line 6: If $f$ is $\hat{\mu}_{s}$-QG and convex, then for any $x\in X$, we have
$\begin{split}f^{*}&\geq f\left(x\right)+\left\langle f^{\prime}\left(x\right),x^{*}-x\right\rangle\geq f\left(x\right)-\left\|f^{\prime}\left(x\right)\right\|\left\|x^{*}-x\right\|\\ &\geq f\left(x\right)-\left\|f^{\prime}\left(x\right)\right\|\sqrt{2\left(f\left(x\right)-f^{*}\right)/\hat{\mu}_{s}},\end{split}\ ,$ (6.3)
which implies that $f^{*}\geq f\left(x\right)-2\left\|f^{\prime}\left(x\right)\right\|^{2}/\hat{\mu}_{s}$. A more refined bound can be derived by using the $\left(\iota_{s-1},\nu_{s-1}\right)$-W-certificate, denote $\mathcal{P}_{s-1}$, from the previous outer iteration. This certificate at $\bar{y}^{s-2}$ can be transformed into a new certificate for point $\bar{y}^{s-1}$. Specifically, if Algorithm 6 enters Line 6 for $s\geq 2$, there exists a point set $\mathcal{P}_{s-1}$ as $\left(\iota_{s-1},\nu_{s-1}\right)$-W-certificate for point $\bar{y}^{s-2}$ in function $f$ by Lemma 6.2. Since $\bar{y}^{s-1}$ is generated in Line 16 of Algorithm 6, it means the $\left\|\bar{y}^{s-1}-\bar{y}^{s-2}\right\|\leq\sqrt{2\theta\tilde{\Delta}_{s-1}/\mu_{s-1}}$, otherwise Algorithm 6 would enter Line 11. By the updating rule, we have $\theta\tilde{\Delta}_{s-1}\leq\Delta_{s-1}$, which implies that $\left\|\bar{y}^{s-1}-\bar{y}^{s-2}\right\|\leq\sqrt{2\Delta_{s-1}/\mu_{s-1}}=\iota_{s-1}$. Furthermore, by Lemma 4.3 indicating $\min_{x\in\mathcal{B}\left(\bar{y}^{s-2};\iota_{s-1}\right)}\psi_{\mathcal{P}_{s-1},f}\left(x\right)\geq f\left(\bar{y}^{s-2}\right)-\Delta_{s-1}$. Hence, applying Lemma 6.1 with $\mathcal{P}=\mathcal{P}_{s-1}$, $\iota=\iota_{s-1}$, $\nu=\nu_{s-1}$, $\bar{y}=\bar{y}^{s-2},$ $\tilde{l}=f\left(\bar{y}^{s-2}\right)-\Delta_{s-1},\tilde{y}=\bar{y}^{s-1}$, and $c=1$ yields that $\tilde{\mathcal{P}}_{s}=\left\{\bar{y}^{s-1}\right\}\cup\mathcal{P}_{s-1}$ is an $\left(\tilde{\iota}_{s},\tilde{\nu}_{s}\right)$-W-certificate for point $\bar{y}^{s-1}$ in function $f$, where $\tilde{\iota}_{s}=2\iota_{s-1}$ and
$\tilde{\nu}_{s}=\frac{1}{2\iota_{s-1}}\left(f\left(\bar{y}^{s-1}\right)-f\left(\bar{y}^{s-2}\right)+3\Delta_{s-1}\right)\ .$ (6.4)
For notation simplicity, we denote $\zeta_{s}=f\left(\bar{y}^{s-1}\right)-f\left(\bar{y}^{s-2}\right)+3\Delta_{s-1}$. By the definition of $\tilde{\nu}_{s}$, we have $\zeta_{s}\geq 0$. Thus, applying Lemma 4.2 with assumption that function $f$ is $\hat{\mu}_{s}$-QG gives
$f\left(\bar{y}^{s-1}\right)-f^{*}\leq\max\left\{\zeta_{s},\frac{1}{2\hat{\mu}_{s}\iota_{s-1}^{2}}\zeta_{s}^{2}\right\}\ .$ (6.5)
Since $\iota_{s-1}^{2}=2\Delta_{s-1}/\mu_{s-1}$, then we have
$2\hat{\mu}_{s}\iota_{s-1}^{2}=\frac{4\hat{\mu}_{s}}{\mu_{s-1}}\Delta_{s-1}\overset{(a)}{\geq}\frac{4\hat{\mu}_{s}}{3\mu_{s-1}}\zeta_{s}\ ,$ (6.6)
where $\left(a\right)$ holds by $f\left(\bar{y}^{s-1}\right)-f\left(\bar{y}^{s-2}\right)\leq 0$. $\zeta_{s}>0$ and (6.6) imply $3\mu_{s-1}\zeta_{s}/4\hat{\mu}_{s}$ is an upper bound of the second term in (6.5). Note that $\hat{\mu}_{s}\leq\mu_{s-1}/2$, hence, we have
$\frac{3\mu_{s-1}}{4\hat{\mu}_{s}}\zeta_{s}\geq\max\left\{\frac{1}{2\hat{\mu}_{s}\iota_{s-1}^{2}}\zeta_{s}^{2},\frac{3}{2}\zeta_{s}\right\}\ .$ (6.7)
Hence, combining (6.7) and (6.5) yields
$\begin{split}f\left(\bar{y}^{s-1}\right)-f^{*}&\leq\max\left\{\zeta_{s},\frac{1}{2\hat{\mu}_{s}\iota_{s-1}^{2}}\zeta_{s}^{2},\frac{3}{2}\zeta_{s}\right\}\leq\frac{3\mu_{s-1}}{4\hat{\mu}_{s}}\zeta_{s}\end{split}\overset{(a)}{\leq}\frac{9\mu_{s-1}}{4\hat{\mu}_{s}}\Delta_{s-1}\ ,$ (6.8)
where $(a)$ holds by $f\left(\bar{y}^{s-1}\right)-f\left(\bar{y}^{s-2}\right)\leq 0$.
Line 11: By Lemma 6.2, Algorithm 6 entering Line 11 means that $\mathcal{AGR}$ generates a $\mathcal{\tilde{P}}_{s}$ as a $\left(\sqrt{2\Delta_{s}/\mu_{s}},\sqrt{\mu_{s}\Delta_{s}/2}\right)$-W-certificate for point $\bar{y}^{s-1}$ in function $f$. Hence, by Lemma 4.2, if $f$ is $\hat{\mu}_{s}$-QG function, then we have
$f\left(\bar{y}^{s-1}\right)-f^{*}\leq\max\left\{\Delta_{s},\frac{2}{\hat{\mu}_{s}}\cdot\frac{1}{2}\mu_{s}\Delta_{s}\right\}\overset{(a)}{=}\Delta_{s}\ ,$ (6.9)
where $(a)$ holds by $\mu_{s}=\hat{\mu}_{s}$.
Lemma 6.4.
Suppose function $f$ is convex, satisfies $\mu^{*}$-QG. In Algorithm 6, if the Quar-Flag returned by $\mathcal{AWG}$ is False, then we have $\hat{\mu}_{s}>\mu^{*}$. Therefore, for any $s\geq 1$, $\hat{\mu}_{s}\geq\mu^{*}/4$ and the number of times Algorithm 6 updates the QG estimate will not exceed $\left\lceil\log_{4}\frac{4\hat{\mu}_{1}}{\mu^{*}}\right\rceil$.
Proof.
The proof is based on contradiction. Suppose $f$ is $\hat{\mu}_{s}$-QG, then it follows from Lemma 6.3 that $\bar{f}-f^{*}\leq\tilde{\Delta}_{s}$. We will show that, under condition $\bar{f}-f^{*}\leq\tilde{\Delta}_{s}$, $\mathcal{AWG}$ can not return Quar-Flag as False.
$\mathcal{AWG}$ is an algorithm that run APEX with level $\tilde{l}=\bar{f}-\left(1+\beta\right)\tilde{\Delta}_{s}$. Hence, it follows from the second result of Proposition 3.1 that
$\omega_{t}\left(f\left(\hat{x}^{t}\right)-\left(\bar{f}-\left(1+\beta\right)\tilde{\Delta}_{s}\right)\right)-\omega_{1}\left(1-\frac{\alpha_{1}}{2}\right)\left(f\left(\hat{x}^{0}\right)-\tilde{l}\right)\leq\bar{L}\left(t\right)\left\|x^{t}-\bar{y}\right\|^{2}\ .$ (6.10)
Since $f\left(\hat{x}^{t}\right)\geq f^{*}$ and $\bar{f}-f^{*}\leq\tilde{\Delta}_{s}$, then we have
$\omega_{t}\beta\tilde{\Delta}_{s}-\omega_{1}\left(1-\frac{\alpha_{1}}{2}\right)\left(f\left(\hat{x}^{0}\right)-\tilde{l}\right)\leq\bar{L}\left(t\right)\left\|x^{t}-\bar{y}\right\|^{2}\ .$ (6.11)
Hence, the above inequality fails implies that $\hat{\mu}_{s}$ is overestimated. On the other hand, by Proposition 4.1, Algorithm 3 will return a W-certificate in at most $m\sqrt{\left(2\iota_{\max}\bar{L}\left(t\right)+6\left(1+\beta\right)\Delta\right)/\left(\beta\Delta\right)}$ gradient calls if $\hat{\mu}_{s}\leq\mu^{*}$. Therefore, once Algorithm 3 returns False, we can obtain $\hat{\mu}_{s}$ is overestimated.
Proposition 6.1.
Let the weighted sequence $\omega_{t}=(t+2)(t+3)/2$, $\alpha_{t}=4/(t+3)$ and $\theta\in(1/2,1)$ in Algorithm 5. If the objective function $f$ is convex, the Algorithm 5 will return at most $O(1)\left(m\sqrt{\left(6\theta+4\theta\bar{L}(t)/\mu\right)/(2\theta-1)}\right)$ gradient calls in total.
Proof.
Algorithm 5 can be viewed as a specific instance of APEX, where the parameters are set as $\tilde{l}=f\left(\bar{y}\right)-\theta\Delta$, $\alpha_{t}=4/(t+3)$, and $\omega_{t}=(t+2)(t+3)/2$. From Proposition 3.1, we have
$\frac{(t+2)(t+3)}{2}\left[f(\hat{x}^{t})-\left(f(\bar{y})-\theta\Delta\right)\right]-3\theta\Delta\leq\bar{L}(t)\left\|x^{t,0}-\bar{y}\right\|^{2}.$
Before Algorithm 5 terminates, it must hold that $f(\hat{x}^{t})\geq\theta\Delta+\underline{f}=f(\bar{y})-(1-\theta)\Delta$. Therefore,
$\frac{(t+2)(t+3)}{2}\,(2\theta-1)\Delta-3\theta\Delta\leq\bar{L}(t)\left\|x^{t,0}-\bar{y}\right\|^{2}.$
It follows that, as soon as $t\geq\sqrt{\left(4\theta\,\bar{L}(t)/\mu+6\theta\right)/(2\theta-1)}$, we must have $\left\|x^{t,0}-\bar{y}\right\|^{2}\geq 2\theta\Delta/\mu$, establishing the claimed result. ∎
Now, we are ready to give the formal statement of convergence result of Algorithm 6.
Theorem 6.1.
Suppose function $f$ is convex, satisfies $\mu^{*}$-QG and is $\left(k,L\right)$-PWS. In Algorithm 6, assume that the number of cuts obeys $m\geq k$, and $\hat{\mu}_{1}\geq\mu^{*}$. Set $\omega_{t}=(t+2)(t+3)/2,\alpha_{t}=4/(t+3)$ in subroutine $\mathcal{AWG}$ and $\mathcal{AGR}$. To find a point set $\mathcal{P}_{s}$ that is an $\left(2\varepsilon/\mu_{s},\varepsilon\right)$-W-certificate for point $\bar{y}^{s-1}$ in function $f$, Algorithm 6 requires a total number of gradient oracle calls bounded by:
$O(1)\cdot m\sqrt{\frac{L}{\mu^{*}}}\left(\log_{4}\frac{4\hat{\mu}_{1}}{\mu^{*}}+\log\frac{\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}}{\varepsilon^{2}}\right)\ .$ (6.12)
Proof.
By Lemma 6.2, when Algorithm 6 enters Line 8 or 11, it produces an $\left(\sqrt{2\Delta_{s}/\mu_{s}},\sqrt{\mu_{s}\Delta_{s}/2}\right)$-W-certificate for point $\bar{y}^{s-1}$ . Hence, to obtain an $\left(2\varepsilon/\mu_{s},\varepsilon\right)$-W-certificate for point $\bar{y}^{s-1}$, it suffices to enforce $\mu_{s}\Delta_{s}\leq 2\varepsilon^{2}$.
By Proposition 4.1 with $\iota_{\max}=\sqrt{\left(2\left(1+\beta\right)\tilde{\Delta}_{s}\right)/\hat{\mu}_{s}}$, together with the third statement in Theorem 5.1 that $\bar{L}\left(t\right)=O(1)L$ if $m\geq k$ , there exist two constants $C_{\beta},C_{\theta}>0$ such that each call to $\mathcal{AWG}$ or $\mathcal{AGR}$ needs at most $C_{\beta}m\sqrt{L/\mu}$ or $C_{\theta}m\sqrt{L/\mu}$ inner iterations, respectively, where $\mu$ is the QG estimate used in $\mathcal{AWG}$ and $\mathcal{AGR}$. Let $\bar{C}=\max\left\{C_{\beta},C_{\theta}\right\}$, then no matter $\mathcal{AWG}$ or $\mathcal{AGR}$, it will return at most $\bar{C}m\sqrt{L/\mu}$ gradient oracle calls.
By Lemma 6.4, the number of quarterings of the QG estimate is bounded by $N_{1}:=\left\lceil\log_{4}\left(\frac{4\hat{\mu}_{1}}{\mu^{*}}\right)\right\rceil$. Let $\tilde{\Delta}^{1,q}$ (resp. $\tilde{\Delta}^{2,q}$) denote $\tilde{\Delta}_{s}$ immediately before (resp. after) the $q$-th quartering of the QG estimates. Since throughout the run we maintain $\hat{\mu}_{s}\geq\mu^{*}/4$, then the total oracle complexity $S(\varepsilon)$ up to achieving $\mu_{s}\Delta_{s}\leq 2\varepsilon^{2}$ can be bounded by
$2\bar{C}m\left(\underbrace{\sqrt{\frac{4L}{\mu^{*}}}N_{1}}_{\text{$\left(\textrm{I}\right)$}}+\underbrace{\sqrt{\frac{L}{\hat{\mu}_{1}}}\log_{1/\theta}\left(\frac{\tilde{\Delta}_{1}}{\tilde{\Delta}^{1,1}}\right)}_{\left(\text{II}\right)}+\underbrace{\sum_{q=1}^{N_{1}-1}\sqrt{\frac{L\cdot 4^{q}}{\hat{\mu}_{1}}}\log_{1/\theta}\left(\frac{\tilde{\Delta}^{2,q}}{\tilde{\Delta}^{1,q+1}}\right)}_{\left(\text{{III}}\right)}+\underbrace{\sqrt{\frac{4L}{\mu^{*}}}\log_{1/\theta}\left(\frac{\mu^{*}\tilde{\Delta}^{2,N_{1}}}{8\varepsilon^{2}}\right)}_{\left(\text{{IV}}\right)}\right).$ (6.13)
Here $\left(\textrm{I}\right)$ accounts for the quarterings QG estimate (Quar-Flag returned by $\mathcal{AWG}$ is False and $f(\hat{x}^{s})\leq\tilde{l}$); $\left(\text{II}\right)$ is the cost from the start to the first quartering; $\left(\text{{III}}\right)$ accumulates the costs between two quarterings; $\left(\text{{IV}}\right)$ is the cost after the last quartering until $\mu_{s}\Delta_{s}\leq 2\varepsilon^{2}$; The factor of 2 arises because the number of oracle calls required by $\mathcal{AWG}$ is less than that required by $\mathcal{AGR}$. For clarity and to provide a unified upper bound, we multiply by 2 in our complexity estimate.
Let $\bar{\Delta}:=\max_{s\geq 1}\left\{\tilde{\Delta}_{s}\right\}$ be the upper bound of $\tilde{\Delta}_{s}$ in the Algorithm 6 running. Before obtaining $\left(2\varepsilon/\mu_{s},\varepsilon\right)$-W-certificate, we have for all $\tilde{\Delta}^{1,q}\geq 2\varepsilon^{2}/\hat{\mu}_{s}\geq 8\varepsilon^{2}/\mu^{*}$. Moreover, $\sum_{q=1}^{N_{1}-1}\sqrt{\frac{L\cdot 4^{q}}{\hat{\mu}_{1}}}\leq\sqrt{\frac{L\cdot 4^{N_{1}}}{\hat{\mu}_{1}}}\leq\sqrt{\frac{16L}{\mu^{*}}}$. Hence, we have
$\left(\text{II}\right)\leq\sqrt{\frac{4L}{\mu^{*}}}\log_{1/\theta}\left(\frac{\mu^{*}\bar{\Delta}}{8\varepsilon^{2}}\right)\ ,\ \left(\text{{III}}\right)\leq\sqrt{\frac{16L}{\mu^{*}}}\log_{1/\theta}\left(\frac{\mu^{*}\bar{\Delta}}{8\varepsilon^{2}}\right)\ ,\ \left(\text{{IV}}\right)\leq\sqrt{\frac{4L}{\mu^{*}}}\log_{1/\theta}\left(\frac{\mu^{*}\bar{\Delta}}{8\varepsilon^{2}}\right).$ (6.14)
Combining (6.13) and (6.14) gives
$S\left(\varepsilon\right)\leq O(1)\cdot m\left(\sqrt{\frac{L}{\mu^{*}}}\log_{4}\left(\frac{4\hat{\mu}_{1}}{\mu^{*}}\right)+\sqrt{\frac{L}{\mu^{*}}}\log_{1/\theta}\left(\frac{\mu^{*}\bar{\Delta}}{\varepsilon^{2}}\right)\right)\ .$ (6.15)
In the following, we give the upper bound for term $\bar{\Delta}$. Algorithm 6 only increases $\tilde{\Delta}_{s}$ at the quartering lines (Line 6), with update rule
$\tilde{\Delta}_{s}\leq\frac{9\mu_{s-1}\Delta_{s-1}}{4\hat{\mu}_{s}}.$ (6.16)
Let $\left\{s_{q}\mid 1\leq q\leq N_{1}\right\}$ be the outer iteration index set where quartering occurs, where subscript $q$ means the $q$-th quartering occurs, and write $\tilde{\Delta}_{s_{q}}$ for the value at exit from iteration $s_{q}$. Since $\Delta_{s_{q}-1}\leq\left(1+\beta\right)\tilde{\Delta}_{s_{q}-1}$ and $\hat{\mu}_{s_{q}}=\mu_{s}$, using (6.16) gives
$\tilde{\Delta}_{s_{q}}\leq\frac{9}{4}\cdot\frac{\mu_{s_{q}-1}}{\mu_{s_{q}}}\Delta_{s_{q}-1}\leq\frac{9\left(1+\beta\right)}{4}\cdot\frac{\mu_{\left(s_{q}\right)-1}}{\mu_{s_{q}}}\tilde{\Delta}_{s_{q}-1}\overset{(a)}{\leq}\frac{9\left(1+\beta\right)}{4}\cdot\frac{\mu_{s_{q-1}}}{\mu_{s_{q}}}\tilde{\Delta}_{s_{q-1}}\ ,$ (6.17)
where the last inequality uses that the QG estimate does not change between the $s_{q-1}$-th outer iteration and the $\left(\left(s_{q}\right)-1\right)$-th outer iteration, i.e., $\mu_{\left(s_{q}\right)-1}=\mu_{s_{q-1}}$. Applying the formula recursively, we have
$\tilde{\Delta}_{s_{q}}\leq\left(\frac{9\left(1+\beta\right)}{4}\right)^{q-1}\frac{\mu_{s_{1}}}{\mu_{s_{q}}}\tilde{\Delta}_{s_{1}}\leq\left(\frac{9\left(1+\beta\right)}{4}\right)^{N_{1}-1}\frac{4\mu_{s_{1}}}{\mu^{*}}\tilde{\Delta}_{s_{1}}\ .$ (6.18)
We now bound $\tilde{\Delta}_{s_{1}}$. If $s_{1}=1$, then
$\tilde{\Delta}_{s_{1}}\leq\frac{2\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}}{\mu_{s_{1}}}$ (6.19)
Otherwise, if $s_{1}>1$, by (6.16) and the initialization bound at stage $s_{1}-1$
$\tilde{\Delta}_{s_{1}}\leq\frac{9\mu_{s_{1}-1}\Delta_{s_{1}-1}}{4\hat{\mu}_{s_{1}}}\overset{(a)}{=}\frac{9\mu_{s_{1}-1}\Delta_{s_{1}-1}}{4\mu_{s_{1}}}\overset{(b)}{\leq}\frac{9\mu_{s_{1}-1}}{4\mu_{s_{1}}}\cdot\frac{2\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}}{\hat{\mu}_{1}}\overset{(c)}{=}\frac{9}{2\mu_{s_{1}}}\cdot\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}\ ,$ (6.20)
where $(a)$ holds by $\hat{\mu}_{s_{1}}=\mu_{s_{1}}$ since $\hat{\mu}_{s_{1}}$ is the QG estimate after exiting the $s_{1}$-th outer iteration, $(b)$ follows from $\Delta_{s_{1}-1}\leq\frac{2\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}}{\hat{\mu}_{1}}$ since the gap always shrinks before the $s_{1}$-th iteration, and $(c)$ holds by $\mu_{s_{1}-1}=\hat{\mu}_{1}$ since before the $s_{1}$-th iteration, the QG estimates do not change. Combining (6.19), (6.20) and (6.18) yields
$\tilde{\Delta}_{s_{q}}\leq\left(\frac{9\left(1+\beta\right)}{4}\right)^{N_{1}-1}\frac{18}{\mu^{*}}\cdot\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}=:\bar{\Delta}\ .$ (6.21)
Substituting (6.21) into (6.15) absorbs the $\beta$ and $\theta$–dependent factors into the hidden $O(1)$, yielding
$O(1)\cdot m\sqrt{\frac{L}{\mu^{*}}}\left(\log_{4}\left(\frac{4\hat{\mu}_{1}}{\mu^{*}}\right)+\log_{1/\theta}\left(\frac{\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}}{\varepsilon^{2}}\right)\right)\ .$ (6.22)
The complexity in Theorem 6.1 is established for an $\left(\varepsilon/\mu_{s},\varepsilon\right)$W-certificate. The W-certificate is verifiable during the algorithmic process, at which stage the W-certificate is guaranteed to be generated correctly. By applying the same proof technique, we obtain the complexity result for the optimality gap, stated in the following Corollary.
Theorem 6.2.
Suppose function $f$ is convex, satisfies $\mu^{*}$-QG and is $\left(k,L\right)$-PWS. In Algorithm 6, the argument that the number of cuts of obeys $m\geq k$ and $\hat{\mu}_{1}\geq\mu^{*}$. To find a solution $\bar{y}^{s-1}$ such that $f\left(\bar{y}^{s-1}\right)-f^{*}\leq\varepsilon$, Algorithm 6 requires a total number of gradient oracle calls bounded by:
$O(1)\cdot m\sqrt{\frac{L}{\mu^{*}}}\left(\log_{4}\frac{4\hat{\mu}_{1}}{\mu^{*}}+\log\frac{\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}}{\mu^{*}\varepsilon}\right)\ .$ (6.23)
Proof.
By Lemma 6.2, Algorithm 6 will generate a point set $\mathcal{P}_{s}$ as $\left(\sqrt{2\Delta_{s}/\mu_{s}},\sqrt{\mu_{s}\Delta_{s}/2}\right)$-W-certificate for point $\bar{y}^{s-1}$ in function $f$. By Lemma 4.2, the W-certificate and $\mu^{*}$-strong convexity implies that $f\left(\bar{y}^{s-1}\right)-f^{*}\leq\max\left\{1,\mu_{s}/\mu^{*}\right\}\Delta_{s}$. To make sure $f\left(\bar{y}^{s-1}\right)-f^{*}\leq\varepsilon$, we need $\mu_{s}\Delta_{s}\leq\min\left\{\mu_{s},\mu^{*}\right\}\varepsilon$. Notice that Theorem 6.1 is to bound the iteration number for condition $\mu_{s}\Delta_{s}\leq 2\varepsilon^{2}$. Hence, replacing $2\varepsilon^{2}$ in the proof procedure of Theorem 6.1 with $\min\left\{\mu_{s},\mu^{*}\right\}\varepsilon$} and noting that ${\mu}_{s}\geq\frac{1}{4}\mu^{*}$, we have the complexity result. ∎
A few remarks regarding this complexity result in Theorem 6.1 and Corollary 6.2 are in order. First, both statements require the assumption $\hat{\mu}_{1}\geq\mu^{*}$. If this assumption fails, then quartering of the strong convexity estimate never occurs, and the resulting complexity reduces to that established in Theorem 5.1 with $\hat{\mu}_{1}$. Second, both results also rely on the condition $m\geq k$. If this assumption fails, the algorithm cannot fully capture the piecewise smooth structure, and the complexity must instead be expressed in terms of $L_{\text{avg}}$, as discussed in Theorem 5.1 and Proposition 3.2. We do not elaborate further here. Third, the oracle complexity in Theorem 6.1 and Corollary 6.2 is dominated by the term $O\left(1\right)\cdot m\sqrt{{L}/{\mu^{*}}}\left(\log{\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}}/{\mu^{*}\varepsilon}\right)$ . This matches the bound in Theorem 5.1, but crucially eliminates the need for prior knowledge of $\mu^{*}$ or the optimal value $f^{*}$. This adaptivity to the true curvature is a significant practical advantage. Fourth, the proposed algorithm is an anytime method, in the sense that it does not require the target accuracy $\varepsilon$ to be specified in advance.
7 Lower oracle complexity bound for $\mu$-QG and $\left(k,L\right)$-PWS optimization
In this section, we establish a lower bound on the oracle complexity for minimizing $\mu$-QG, $\left(k,L\right)$-PWS function. Our main result (Theorem 7.2) shows that every deterministic first-order method requires at least $\Omega\left(k\sqrt{{L}/{\mu}}\log\left({1}/{\varepsilon}\right)\right)$ oracle calls to reach an $\varepsilon$-optimal solution. When the number of cuts $m$ is set to $k$, this matches exactly the upper bound guarantees proved in Section 5 and Section 6. Hence, thereby certifying that Algorithm 4 and 6 are optimal up to constant factors.
The proof follows the two-stage approach pioneered by [11, 12]. First, we show that the lower complexity of any first-order deterministic algorithm (Definition 6) is lower bounded by that of any zero-respecting algorithm (Definition 7) in Proposition 7.1. It thus suffices to prove the lower bound within the zero-respecting algorithm class. Second, inspired by Nesterov’s construction [65], we exhibit a simple $\mu$-strongly convex, $\left(k,L\right)$-PWS function for which every zero-respecting method indeed incurs $\Omega\left(k\sqrt{{L}/{\mu}}\log\left({1}/{\varepsilon}\right)\right)$ queries. Since strongly convex functions are also $\mu$-QG functions, this result also applies to $\mu$-QG functions.
We begin by introducing precise definitions under consideration, including function class (Definition 4), orthogonal connection matrix (Definition 5), first-order deterministic algorithm (Definition 6) and first-order zero-respecting sequences and algorithm (Definition 7).
We index our function families by their input dimension, specifically:
Definition 4.
We denote by $\mathcal{F}^{d}$ a class of $d$-dimensional functions; that is, each $f\in\mathcal{F}^{d}$ is a mapping $f:\mathbb{R}^{d}\to\mathbb{R}$.
Two function classes may be related by transformations such as lifting or compression. If there exists an orthogonal matrix $U$ that maps every function in one class to a corresponding function in the other, then $U$ is called a connection matrix between those two function classes. Moreover, if matrix $U$ also satisfies the orthogonality condition, then $U$ is referred to as an orthogonal connection matrix.
Definition 5 (Orthogonal connection matrix).
We say a matrix $U\in\mathbb{R}^{\tilde{d}\times d}$ is orthogonal connection matrix from function class $\mathcal{F}^{d}$ to function class $\mathcal{F}^{\tilde{d}}$ if $U^{\top}U=I_{d}$ and for every function $f\in\mathcal{F}^{d}$, there exist a function $f_{U}\left(x\right)=f\left(U^{\top}x\right)\in\mathcal{F}^{\tilde{d}}$.
The preceding definitions separate the single-dimensional construction of Carmon et al [11, 12] into two input dimensions, $\tilde{d}$ and $d$. However, not every pair of function classes $\mathcal{F}^{d}$ and $\mathcal{F}^{\tilde{d}}$ admits an orthogonal connection: for example, there is no orthogonal matrix mapping general quadratic functions in $\mathcal{F}^{d}$ onto cubic functions in $\mathcal{F}^{\tilde{d}}$. Leveraging an orthogonal connection matrix, one can transfer lower-bound results between these classes by invoking the orthogonal invariance principle of Nemirovski and Yudin [61].
Next, the two classes of algorithms under consideration are defined as follows.
Definition 6 (First-order deterministic algorithm).
We say algorithm class $\mathbb{A}_{\text{det}}$ is the first-order deterministic algorithm class if for all algorithm $\mathcal{A}\in\mathbb{A}_{\text{det}}$ on operating on function $f:\mathbb{R}^{d}\to\mathbb{R}$, the producing iterates satisfy:
$x^{i}=\mathcal{A}\left(\nabla f\left(x^{1}\right),\ldots,\nabla f\left(x^{i-1}\right)\right)\ ,\ \forall i\in\mathbb{N}\ .$ (7.1)
Definition 7 (First-order zero-respecting algorithm).
For any $f:\mathbb{R}^{d}\to\mathbb{R}$, we say that sequence $\left\{x^{1},x^{2},\ldots\right\}$ is a first-order zero-respecting sequence with respect to $f$ if
$\operatorname{supp}\left\{x^{t}\right\}\subseteq\cup_{s<t}\operatorname{supp}\left\{\nabla f\left(x^{s}\right)\right\}\ ,\ \text{for each }t\in\mathbb{N}.$ (7.2)
We say algorithm class $\mathbb{A}_{\text{zr}}$ is a first-order zero-respecting algorithm class if any algorithm $\mathcal{A}\in\mathbb{A}_{\text{zr}}$ is a first-order zero-respecting algorithm, i.e., for any $f:\mathbb{R}^{d}\to\mathbb{R}$, the iterate sequence $\mathcal{A}\left[f\right]$ is first-order zero-respecting with respect to $f$.
Zero-respecting algorithms were first introduced by Carmon et al [11, 12], who also developed a generalized $p$-th order framework. In this work, we restrict our attention to the first-order case and accordingly present only the first-order zero-respecting algorithm. Informally, a first-order algorithm is said to be zero-respecting if, at every iteration, it refrains from querying any coordinate whose partial derivative has remained identically zero throughout all previous iterations. When initialized at the origin, most standard first-order methods—such as gradient descent and conjugate gradient—naturally satisfy this property.
Having fixed the function family $\mathcal{F}^{d}$ and the algorithm class $\mathbb{A}$, we now define the worst-case iteration complexity required to attain an $\varepsilon$-optimal solution. Denote any sequence $\left\{x^{t}\right\}_{t\in\mathbb{N}}$ generated by algorithm $\mathcal{A}$ on function $f$ as $\mathcal{A}\left[f\right]$, define its $\varepsilon$-complexity on $f:\mathbb{R}^{d}\to\mathbb{R}$ by
$T_{\varepsilon}\left(\mathcal{A}\left[f\right],f\right):=\inf\left\{t\in\mathbb{N}:f\left(x^{t}\right)-f^{*}\leq\varepsilon\right\}\ .$ (7.3)
With this setup, we define the complexity of algorithm class $\mathbb{A}$ on function $\mathcal{F}^{d}$ as
$\mathcal{T}_{\varepsilon}\left(\mathbb{A},\mathcal{F}^{d}\right):=\inf_{\mathcal{A}\in\mathbb{A}}\sup_{f\in\mathcal{F}^{d}}T_{\varepsilon}\left(\mathcal{A}\left[f\right],f\right)\ .$ (7.4)
This quantity captures the minimal number of iterations that the best algorithm in $\mathbb{A}$ must make to guarantee $f\left(x^{t}\right)-f^{*}\leq\varepsilon$ for every $f\in\mathcal{F}^{d}$.
Having established the necessary definitions, we now state Proposition 7.1, which provides a pivotal reduction and clarifies the relationship between the class of deterministic first-order algorithms and the zero-respecting algorithm class.
Proposition 7.1.
Suppose there exist a matrix $U\in\mathbb{R}^{\tilde{d}\times d}$ that is orthogonal connection matrix from function class $\mathcal{F}^{d}$ to function class $\mathcal{F}^{\tilde{d}}$ with $\tilde{d}>2d$ and $\varepsilon>0$. Then $\text{$\mathcal{T}$}_{\varepsilon}\left(\mathbb{A}_{\text{det}},\mathcal{F}^{\tilde{d}}\right)\geq\mathcal{T}_{\varepsilon}\left(\mathbb{A}_{\text{zr}},\mathcal{F}^{d}\right)$.
By Proposition 7.1 , establishing a lower bound for any deterministic first-order method on the $\tilde{d}-$dimensional class $\mathcal{F}^{\tilde{d}}$ reduces to proving a corresponding bound for zero-respecting methods on the $d$-dimensional class $\mathcal{F}^{d}$. Accordingly, we henceforth focus on deriving the worst-case complexity of the zero-respecting class $\mathbb{A}_{\text{zr}}$ over $\mathcal{F}^{d}$.
The hard instance for $(k,L)$-PWS function is adapted from the “chain-like” quadratic function presented in Nesterov’s introductory textbook [65]. Specifically, the “chain-like” quadratic function is
$\phi_{\mu,L}\left(x\right)=\frac{\mu}{2}\left\|x\right\|^{2}+\frac{L-\mu}{8}x^{\top}Ax-\frac{L-\mu}{4}e_{1}^{\top}x\ ,$ (7.5)
where
$A=\left[\begin{array}[]{ccccc}2&-1\\ -1&2&-1\\ &-1&\ddots&\ddots\\ &&\ddots&2&-1\\ &&&-1&2\end{array}\right]\ .$
The important properties of $\phi_{\mu,L}\left(x\right)$ are that
  1. 1.
    $\phi_{\mu,L}\left(x\right)$ is $\mu$-strongly convex and $L$-smooth,
  2. 2.
    $\nabla\phi_{\mu,L}\left(0\right)=-\frac{L-\mu}{4}e_{1}$,
  3. 3.
    $\nabla\phi_{\mu,L}\left(x\right)=\mu x+\frac{L-\mu}{4}Ax-\frac{L-\mu}{4}e_{1}$ and the special tri-diagonal structure of matrix $A$ such that if $x\in\operatorname{Span}\left\{e_{i_{1}},e_{i_{2}},\ldots,e_{i_{j}}\right\}$ then $\nabla\phi_{\mu,L}\left(x\right)\in\operatorname{Span}\left\{e_{i_{1}},e_{i_{2}},\ldots,e_{i_{j}},e_{i_{j}+1}\right\}$,
  4. 4.
    $x^{*}=\operatornamewithlimits{argmin}_{x\in\mathbb{R}^{d}}\phi_{\mu,L}\left(x\right)$ satisfies $x_{(i)}^{*}=\left(\frac{\sqrt{L/\mu}-1}{\sqrt{L/\mu}+1}\right)^{i},i=1,\ldots$, where $x_{(i)}^{*}$ is $i$-th element of variable $x^{*}$.
This quadratic example is known to achieve the worst-case $\Omega\left(\sqrt{\frac{L}{\mu}}\log\left(\frac{1}{\varepsilon}\right)\right)$ iteration complexity for first-order methods on $\mu$-strongly convex, $L$-smooth functions. To extend this to the $\left(k,L\right)$-piecewise smooth setting, partition $\mathbb{R}^{d}$ into $k$ blocks of dimensions $n_{1},\ldots,n_{k}$(with $\sum_{i=1}^{k}n_{i}=d$) and define
$\phi_{\mu,L}^{k}\left(x\right)=\max_{i\in\left[k\right]}\left\{\phi_{\mu,L}\left(x_{(i)}\right)\right\}\ ,\ x=\left[x_{(1,:)}^{\top},x_{(2,:)}^{\top},\ldots,x_{(k,:)}^{\top}\right]\in\mathbb{R}^{d}\ ,\ x_{(i,:)}\in\mathbb{R}^{n_{i}}\ ,\ \forall i\in\left[k\right]\ .$ (7.6)
In Theorem 7.2, we show that every first-order zero-respecting algorithm in $\mathbb{A}_{\text{zr}}$ requires $\Omega\left(k\sqrt{\frac{L}{\mu}}\log\left(\frac{1}{\varepsilon}\right)\right)$ iterations to achieve an $\varepsilon$-accurate solution on $\phi_{\mu,L}^{k}\left(x\right)$, thereby establishing the desired lower bound.
Proposition 7.2.
For any $x^{0}\in\mathbb{R}^{\infty}$ and any constants $L>\mu>0$, there exist a $\mu$-strongly convex and $\left(k,L\right)-\text{PWS}$ function $f(x)$ for any first-order deterministic $\mathcal{A}_{\det}\in\mathbb{A}_{\det}$, we have
$\left\|x^{t}-x^{*}\right\|^{2}$ $\geq\left(\frac{\sqrt{L/\mu}-1}{\sqrt{L/\mu}+1}\right)^{2t/k}\left\|x^{0}-x^{*}\right\|^{2}\ ,$ (7.7)
$f\left(x^{t}\right)-f\left(x^{*}\right)$ $\geq\frac{\mu}{2}\left(\frac{\sqrt{L/\mu}-1}{\sqrt{L/\mu}+1}\right)^{2t/k}\left\|x^{0}-x^{*}\right\|^{2}\ ,$ (7.8)
where $x^{*}$ is the unique unconstrained minimum of function $f$.
Proof.
It suffices to prove that the claims (7.7) and (7.8) hold for any first-order zero-respecting algorithm $\mathcal{A}_{zr}$ by Proposition 7.1. We test $\mathcal{A}_{zr}$ on the PWS instance $f\left(x\right)=\phi_{\mu,L}^{k}\left(x\right)$, where in the definition (7.6) we set $n_{i}=\infty,i=1,\ldots,k$. For notation simplicity, we denote $q=\frac{\sqrt{L/\mu}-1}{\sqrt{L/\mu}+1}$. Hence, by the fourth property (Item 4) of $\phi_{\mu,L}\left(x\right)$ that $x^{*}=\operatornamewithlimits{argmin}_{x\in\mathbb{R}^{\infty}}f\left(x\right)$ has a block-coordinate structure that is $x^{*}=\left[\left(x_{(1,:)}^{*}\right)^{\top},\left(x_{(2,:)}^{*}\right)^{\top},\ldots,\left(x_{(k,:)}^{*}\right)^{\top}\right]$, where $x_{(j,i)}^{*}=q^{i}$. Without loss of generality, we assume algorithm starts from $x^{0}=0$. Then by construction, we have
$\left\|x^{0}-x^{*}\right\|^{2}=\sum_{j=1}^{k}\sum_{i=1}^{n_{j}}x_{(j,i)}^{*}=\frac{kq^{2}}{1-q^{2}}\ .$ (7.9)
By the third property (Item 3) of function $\phi_{\mu,L}\left(x\right)$ and $\mathcal{A}_{zr}$ is a zero-respecting algorithm that $\operatorname{supp}\left\{x^{t}\right\}\subseteq\cup_{s<t}\operatorname{supp}\left\{\nabla f\left(x^{s}\right)\right\}$. Hence, if the total iteration budget $t$ is split into $t_{1},\ldots,t_{k}$ steps devoted to each of the $k$ blocks, then in block $i$ at most the first $t_{i}$ coordinates of $x_{(i)}^{t}$ can be nonzero. Consequently, the error satisfies
$\begin{split}&\left\|x^{t}-x^{*}\right\|^{2}\geq\sum_{i=1}^{k}\left\|x_{(i,:)}^{t}-x_{(i,:)}^{*}\right\|^{2}\overset{(a)}{\geq}\sum_{i=1}^{k}\left\|x_{(i,t_{i}+1:\infty)}^{*}\right\|^{2}\\ &\overset{(b)}{=}\sum_{i=1}^{k}\frac{q^{2(t_{i}+1)}}{1-q^{2}}\overset{(c)}{=}\left\|x^{0}-x^{*}\right\|^{2}\sum_{i=1}^{k}\frac{q^{2t_{i}}}{k}\overset{(d)}{\geq}\left\|x^{0}-x^{*}\right\|^{2}q^{2t/k}\end{split}\ ,$ (7.10)
In the above, $(a)$ holds by discarding the first difference norm between the first $t_{i}$ elements of $x_{(i,:)}^{t}$ and $x_{(i,:)}^{*}$ , $(b)$ follows from $x_{(j,i)}^{*}=q^{i}$, $(c)$ holds by (7.9) and $(d)$ is because $g:t\to q^{2t}$ is convex, then $\frac{1}{k}\sum_{i=1}^{k}g\left(t_{i}\right)\geq g\left(\frac{1}{k}\sum_{i=1}^{k}t_{i}\right)=q^{2\bar{t}}$. Hence, we complete the first result (7.7). Combining it and $f\left(x\right)$ is $\mu$-strongly convex, we complete the proof of the second result (7.8). ∎
8 Numerical Study
In this section, we evaluate the performance of our optimization algorithms on two tasks, one is minimize a finite max of convex quadratic functions (Section 8.1) and the other is two stage stochastic linear programming (Section 8.2). All experiments were conducted on a Mac mini M2 Pro with 32 GB of RAM.
8.1 Finite max of convex quadratic functions (MAXQUAD)
We focus on the following optimization problem:
$\min_{x\in\mathbb{R}^{d}}\max_{i\in[k]}\ \ \frac{1}{2}x^{\top}A_{i}x+b_{i}^{\top}x+c_{i}\ ,$ (MAXQUAD)
where matrix $A_{i}\in\mathbb{R}^{d\times d}$ is a randomly generated positive definite matrix, and $\left\{b_{i},c_{i}\right\}_{i=1}^{k}$ are drawn from normal distribution randomly. The piece number can be determined manually by choosing $k$. Furthermore, we can control the QG modulus (strong convexity in MAXQUAD) by setting the eigenvalues of matrix $A_{i}$. For small-scale instances ($d=300,500$), each matrix $A_{i}$ is constructed with eigenvalues following an arithmetic sequence from $\mu$ to $L$ and distinct, randomly generated eigenvectors. For large-scale problems (d=10,000d=10{,}000 or d=20,000d=20{,}000), to save memory, all $A_{i}$ share a single set of eigenvectors, while the eigenvalues are generated independently for each matrix.
The performance of rAPEX on MAXQUAD.
Experiment setting. We first evaluate the performance of rAPEX on small-scale MAXQUAD with varying piece number $k$, smoothness $L$ and initial QG estimates $\hat{\mu}_{1}$, as reported in Table 2. In all experiments, we set the dimension ($d$) to $300$, cut number ($m$) to $50$, the optimal QG modulus ($\mu^{*}$) to $1.0$, and terminate rAPEX once rAPEX improves the upper bound (Algorithm 6 enters Line 16) with $\Delta_{s}\leq 10^{-6}$. Ipopt [71] is used as a baseline to provide optimal objective $(f^{*})$.
Key observations. Five observations can be drawn from Table 2. First, the choice of initial $\hat{\mu}_{1}$ is robust. Across $\hat{\mu}_{1}$ in $\left\{1,10,100,1000\right\}$, the final objective values are nearly identical to the baseline’s result. Second, the algorithm does not always recover the true $\mu^{*}$. Even when the final estimated QG modulus $\hat{\mu}_{s}$ exceeds $\mu^{*}$, the objective value remains close to the optimum (absolute error is almost at $10^{-6}$ order). For instance, when $k=50$, $L=10.0$ and $\hat{\mu}_{1}=100$, the final QG estimate is $\hat{\mu}_{s}=25$, much larger than $\mu^{*}=1.0$, yet the absolute error of the objective is small. Third, the iteration trends align with our intuition about how the initial QG estimate affects convergence. A larger initial $\hat{\mu}_{1}$ leads to more aggressive lower bound updates, which generally results in fewer iterations when no adjustment to the estimate is needed. This is clearly demonstrated when $(k,L)=(200,1000)$, where the final QG estimate $\hat{\mu}_{s}$ remains equal to the initial $\hat{\mu}_{1}$ throughout optimization. In these cases, larger initial values consistently reduce the iteration count through more aggressive updates. However, for the rest cases, while this trend holds for moderate initial estimates ($\hat{\mu}_{1}\in\left\{1,10\right\}$), it breaks down at $\hat{\mu}_{1}\in\left\{100,1000\right\}$. With such an aggressive initial estimate, the algorithm must spend additional iterations adjusting $\hat{\mu}_{1}$ downward during optimization, ultimately leading to slower convergence. Fourth, the negative error of the objective value means that the objective value of rAPEX is better than the baseline. Finally, as highlighted in Lemma 3.2 and the corresponding remark, the convergence rate’s dependence on the smoothness parameter is governed by the specific pieces traversed during the iterations. Thus, even when the number of cuts is less than the total number of pieces, rAPEX can still effectively capture the underlying PWS structure in many cases—such as $(k,L)=(200,1000)$ and (k,L)=(50,100)(k,L)=(50,100)—and have good convergence performance. In these instances, having fewer cuts than pieces does not hamper the overall convergence behavior. However, for $(k,L)=(200,5.0)$ and $(k,L)=(200,10.0)$, we observe that significantly more iterations are required. We suspect this is because the gap between $\mu$ and $L$ is small, resulting in many more active local pieces, and consequently, $m=50$ cuts are insufficient to capture the local PWS structure.
Table 2: Comparison of rAPEX on MAXQUAD in different initial QG modulus $\hat{\mu}_{1}$
[…]
50 5.00E+$0$ 1 1.00E+$0$ 4043 -6.91E-01 -6.91E-01 2.08E-07
10 1.00E+01 2218 -6.91E-01 -6.91E-01 -4.26E-08
100 6.25E+$0$ 2494 -6.91E-01 -6.91E-01 1.36E-07
1000 3.91E+$0$ 2542 -6.91E-01 -6.91E-01 6.06E-08
1.00E+01 1 1.00E+$0$ 2785 -1.74E-01 -1.74E-01 2.09E-07
10 1.00E+01 2003 -1.74E-01 -1.74E-01 1.19E-07
100 2.50E+01 1565 -1.74E-01 -1.74E-01 3.77E-07
1000 1.56E+01 1903 -1.74E-01 -1.74E-01 1.02E-07
1.00E+02 1 1.00E+$0$ 978 1.22E+$0$ 1.22E+$0$ 2.68E-07
10 1.00E+01 753 1.22E+$0$ 1.22E+$0$ 1.82E-07
100 1.00E+02 562 1.22E+$0$ 1.22E+$0$ 2.45E-07
1000 6.25E+01 671 1.22E+$0$ 1.22E+$0$ 3.83E-07
1.00E+03 1 1.00E+$0$ 1152 2.34E+$0$ 2.34E+$0$ 2.25E-07
10 1.00E+01 810 2.34E+$0$ 2.34E+$0$ 1.92E-07
100 1.00E+02 482 2.34E+$0$ 2.34E+$0$ 1.94E-07
1000 2.50E+02 350 2.34E+$0$ 2.34E+$0$ 3.26E-07
200 5.00E+$0$ 1 1.00E+$0$ 195088 7.41E-01 7.41E-01 1.10E-05
10 1.00E+01 312249 7.41E-01 7.41E-01 8.04E-07
100 1.00E+02 136814 7.41E-01 7.41E-01 1.08E-06
1000 1.00E+03 128509 7.41E-01 7.41E-01 2.02E-06
1.00E+01 1 1.00E+$0$ 408943 1.06E+$0$ 1.06E+$0$ 3.55E-07
10 1.00E+01 33833 1.06E+$0$ 1.06E+$0$ 5.57E-07
100 2.50E+01 32431 1.06E+$0$ 1.06E+$0$ 4.49E-07
1000 1.56E+01 15204 1.06E+$0$ 1.06E+$0$ 5.01E-07
1.00E+02 1 1.00E+$0$ 1192 2.22E+$0$ 2.22E+$0$ 3.26E-07
10 1.00E+01 984 2.22E+$0$ 2.22E+$0$ 2.78E-07
100 1.00E+02 684 2.22E+$0$ 2.22E+$0$ 1.53E-07
1000 6.25E+01 761 2.22E+$0$ 2.22E+$0$ 4.02E-07
1.00E+03 1 1.00E+$0$ 1080 2.96E+$0$ 2.96E+$0$ 1.56E-07
10 1.00E+01 679 2.96E+$0$ 2.96E+$0$ 1.65E-07
100 1.00E+02 329 2.96E+$0$ 2.96E+$0$ 2.22E-07
1000 1.00E+03 190 2.96E+$0$ 2.96E+$0$ 3.33E-07
We further plot the convergence of the upper bound $(\bar{f})$ and lower bound $(\underline{f})$ in Figure 6 and Figure 6, where the $y$-axis denotes the absolute error $\bar{f}-f^{*}$ or $\underline{f}-f^{*}$. Four observations follow. First, excluding the $\hat{\mu}_{1}=1000$ case in Figure 6, convergence is generally faster with larger initial $\hat{\mu}_{1}$, since a larger estimate implies a more aggressive lower bound update, which indirectly accelerates level adjustment and upper bound improvement. Second, inaccurate QG estimates can lead to incorrect lower bounds exceeding the optimal value, resulting in vertical segments in Figure 6 where rAPEX adjusts the QG and renews the lower bound estimates. Third, linear convergence is evident in both datasets, which matches the theoretical result in Theorem 6.1 and Corollary 6.2. In Figure 6, the total number of pieces $(50)$ matches the cut number of rAPEX setting, while in the Figure 6, the cut number $(50)$ is smaller than the total PWS pieces $(200)$. Nevertheless, since rAPEX relies on local pieces during iterations, it successfully captures the PWS structure and still demonstrates linear convergence. Finally, the lower bound and upper bound improve is almost symmetric, which is consistent that our certificate implies a constant factor of the optimality gap.
Figure 5: Absolute error of upper bound and lower bound of rAPEX v.s. Iteration on MAXQUAD with $k=50,L=10$ in Table 2.
Figure 6: Absolute error of upper bound and lower bound of rAPEX v.s. Iteration on MAXQUAD with $k=200,L=1000$ in Table 2.
Comparison with other methods on MAXQUAD.
We compare the following methods:
Experimental setup. For all experiments summarized in Table 3, we fix the number of pieces at $k=50$. For $n=500$, we set the strong convexity parameter $\mu=1.0$ and vary the smoothness parameter $L$ over the set $\left\{5,10,100,1000\right\}$. For large-scale instances, we do not explicitly control $\mu$ or $L$; instead, eigenvalues are sampled randomly and a shared set of eigenvectors is used for all matrices to reduce memory usage.
Method-specific settings. We impose $\hat{\mu}_{1}=100$ for rAPEX and a box constraint $10^{3}\leq x\leq 10^{3}$ for rAPL, which requires optimization over a bounded set. Since both SRD and SRD_adapt compute the gradient of every function in the max structure at each iteration, we report their total number of gradient oracle calls as the product of the number of iterations and the number of pieces ($k=50$), enabling a fair comparison. Across all experiments, we impose a maximum time limit of 3 hours and a cap of 20,00020,\!000 iterations.
Performance criteria. For small-scale problems ($n=500$), we use the Ipopt solution as the ground-truth baseline; for large-scale problems (n=10,000n=10{,}000 and n=20,000n=20{,}000), the baseline is set as the minimum value achieved by all tested methods. Table 3 reports the number of gradient oracle calls required by each method to reach $f(x)-f^{*}\leq 10^{-6}$ (small-scale) and $f(x)-f^{*}\leq 10^{-5}$ (large-scale).
Key findings. We highlight two main observations from Table 3. First, for small-scale problems, rAPEX consistently requires fewer gradient oracle calls than SRD, SRD_adapt, and rAPL across all tested smoothness regimes. The proximal bundle method achieves the lowest number of oracle calls only when $L/\mu=100/1.0$ and $1000/1.0$, but in all other settings, rAPEX is competitive or superior. Since we set the number of cuts $m=50$, rAPEX can fully exploit the piecewise smoothness, yielding strong empirical performance in line with theoretical expectations. Second, for large-scale problems, rAPEX maintains a clear advantage, requiring dramatically fewer gradient oracle calls than SRD and SRD_adapt. Competing approaches such as rAPL and the proximal bundle method are unable to handle these larger problem sizes, as indicated by the dashes in the table. Third, we observe an interesting phenomenon with the proximal bundle method: it fails to achieve the required accuracy for $L/\mu=5/1.0$ and $10/1.0$ within the given time and iteration limits, but outperforms other methods when $L/\mu=100/1.0$ and $1000/1.0$. We suspect that the proximal bundle method cannot efficiently exploit the piecewise-smooth (PWS) structure of the problem. This potential limitation seems to be more pronounced when $L/\mu$ is small, possibly because the number of active local pieces increases in such regimes. Evidence supporting this suspicion can be seen in the increasing number of gradient calls required by rAPEX as $L/\mu$ decreases (for example, in the cases with $(k,L)\in\left\{(200,5),(200,10)\right\}$ in Table 2).
Table 3: Comparison of gradient oracle calls for different methods on MAXQUAD with $k=50$. Each cell reports the mean (top) and standard deviation (bottom) over 5 runs with different random seeds.
$d$ $L/\mu$ rAPEX rAPL prox-bundle SRD SRD_adapt
500 5/1.0 2.40E+03 3.28E+03 - 2.34E+04 2.16E+04
$\pm\textbf{9.48E+01}$ $\pm$ 8.58E+02 $\pm$ 2.48E+03 $\pm$ 1.58E+03
10/1.0 2.22E+03 3.34E+03 - 1.58E+05 1.81E+05
$\pm$ 9.46E+01 $\pm$ 4.07E+02 $\pm$ 7.44E+03 $\pm$ 7.40E+03
100/1.0 7.40E+02 1.25E+03 4.69E+02 4.28E+04 4.33E+04
$\pm$ 1.51E+02 $\pm$ 7.07E+01 $\pm$ 5.43E+01 $\pm$ 9.16E+03 $\pm$ 8.93E+03
1000/1.0 2.00E+02 9.05E+02 1.02E+02 1.17E+06 1.33E+06
$\pm$ 8.08E+$0$ $\pm$ 4.92E+01 $\pm$ 3.20E+$0$ $\pm$ 5.24E+05 $\pm$ 5.63E+05
10000 - 2.49E+03 - - 2.69E+04 2.59E+04
$\pm$ 1.66E+02 $\pm$ 1.04E+03 $\pm$ 6.35E+02
20000 - 2.70E+03 - - 2.77E+04 2.48E+04
$\pm$ 1.24E+02 $\pm$ 5.08E+02 $\pm$ 8.37E+02
8.2 Two-stage stochastic linear programming
In this subsection, we consider the two-stage stochastic linear programming:
$\min_{l_{x}\leq x\leq u_{x},x\in\mathbb{R}^{n_{1}}}\ c_{1}^{\top}x+\mathbb{E}_{\xi\sim\Xi}\left[Q\left(x;\xi\right)\right]\ \operatornamewithlimits{s.t.}\ A_{1}x?b_{1}\ ,\ A_{1}\in\mathbb{R}^{m_{1}\times n_{1}}$ (8.1)
where
$\begin{aligned} Q\left(x;\xi\right)=\min_{l_{y}\leq y\leq u_{y},y\in\mathbb{R}^{n_{2}}}&c\left(\xi\right)^{\top}y\ \\ \operatornamewithlimits{s.t.}\ &A_{1}\left(\xi\right)x+A_{2}\left(\xi\right)y?b_{2}\left(\xi\right),A_{1}(\xi)\in\mathbb{R}^{m_{2}\times n_{1}},A_{2}(\xi)\in\mathbb{R}^{m_{2}\times n_{2}}\end{aligned}.$ (8.2)
Here, we use “?” to denote the possible $\geq"$, ="=" and $\leq"$.
We use benchmark datasets from [53], which are available at website  [^2] https://pages.cs.wisc.edu/~swright/stochastic/sampling/. We consider different sample size of the stochastic scenario, the specific problem dimension is described in the following Table 4.
Table 4: Problem dimensions and number of scenarios for two-stage stochastic linear programming test instances.
Problem $n_{1}$ $m_{1}$ $n_{2}$ $m_{2}$ Total. Scen ($|\Xi|$)
storm 121 185 1259 528 {8,27,125,1000}
term20 63 3 764 124 {50,300}
For the two-stage stochastic linear programming experiments, we compare the performance of rAPEX and rAPL with the number of cutting planes set to $m=50$. We set the initial $\hat{\mu}_{1}=100$ for rAPEX and enforce box constraints in rAPL, setting the lower and upper bounds to $10^{3}$ and $10^{3}$, respectively. All subproblems for both algorithms are solved using Gurobi [32]. Table 5 summarizes the number of gradient oracle calls required by each method to reach an optimality gap of $f(x)-f^{*}\leq\left\{10^{-3},10^{-4}\right\}$. The reference value $f^{*}$ is computed by solving the corresponding deterministic linear program—incorporating all scenarios as constraints—using Gurobi. Two key observations can be drawn from the results. First, rAPEX consistently outperforms rAPL across all datasets and tolerance levels, with the sole exception of the term20 instance with 300 scenarios at the $10^{-3}$ tolerance, where rAPL performs slightly better. Second, as the required tolerance becomes more stringent, the advantage of rAPEX becomes even more pronounced: rAPL often fails to find a satisfactory solution within the 3-hour time limit and 20,000-iteration cap, especially for a $10^{-4}$ tolerance. In contrast, the number of iterations required by rAPEX to reach $10^{-4}$ accuracy remains nearly the same as that needed for $10^{-3}$ accuracy. For example, in the term20 instance with 50 scenarios, the optimality gap decreases directly below $10^{-4}$, resulting in an identical iteration count for both tolerance levels.
Table 5: Comparsion of gradient oracle calls for two methods on two-stage stochastic linear programming
dataset $f^{*}$ $10^{-3}$ $10^{-4}$
rAPEX rAPL rAPEX rAPL
storm8 1.500384E+07 282 446 332 -
storm27 1.500192E+07 510 620 527 -
storm125 1.500151E+07 398 704 452 -
storm1000 1.522862E+07 384 555 2448 -
term20(50) 2.559739E+05 1254 1258 1254 1696
term20(300) 2.549250E+05 1341 1308 1578 1747
Appendix
Appendix A Proof of Proposition 7.1
We first give some basic notations and definitions for the proof of Proposition 7.1. Since these notations and definitions are only used for lower complexity proof, we define it here for readers convenience. The whole proof idea is follows from [11, 12]. Compared with these two work, we focus more on optimality criteria $f\left(x\right)-f^{*}\leq\varepsilon$ and we differ the function class that $\mathcal{F}$ in two different dimensions $d$ and $\tilde{d}$.
For $v\in\mathbb{R}^{d}$, we let $\operatorname{supp}\left\{v\right\}:=\left\{i\in\left[d\right]:v_{i}\neq 0\right\}$ denote the support (non-zero indices) of $v$. For algorithm $\mathcal{A}$ and function $f$, we denote the iterates generated by algorithm $\mathcal{A}$ on function $f$ as $\mathcal{A}\left[f\right]$.
The core of the proofs of Proposition 7.1 is based on the following Lemmas.
Lemma A.1.
Let algorithm $\mathcal{A}\in\mathbb{A}_{\text{det}}$. For any function $f:\mathbb{R}^{d}\to\mathbb{R}$ in $\mathcal{F}$, there exists an orthogonal matrix $\in\mathbb{R}^{\tilde{d}\times d}, UU$$^{\top}U=I_{d}$ and an algorithm $\mathcal{Z}_{\mathcal{A}}\in\mathbb{A}_{\text{zr}}$ such that $f_{U}\left(x\right):=f\left(U^{\top}x\right)$ satisfies that the first $d$ iterates in sequence $\mathcal{Z}_{\mathcal{A}}\left[f\right]$ and $U^{\top}\mathcal{A}\left[f_{U}\right]$ are identical. Furthermore, for any $\varepsilon>0$,
$T_{\varepsilon}\left(\mathcal{A},f_{U}\right)>d\ \text{or\ }T_{\varepsilon}\left(\mathcal{A},f_{U}\right)=T_{\varepsilon}\left(\mathcal{Z}_{\mathcal{A}},f\right)\ .$ (A.1)
Proof.
For notation simplicity, we define the sequence that zero-respecting algorithm $\mathcal{Z}_{\mathcal{A}}$ on function $f$ is $\mathcal{Z}_{\mathcal{A}}\left[f\right]=\left\{z^{1},\ldots,\right\}$ and the sequence that Algorithm $\mathcal{A}$ on function $f_{U}:\mathbb{R}^{\tilde{d}}\to\mathbb{R}$ is $\mathcal{A}\left[f_{U}\right]=\left\{a^{1},\ldots,\right\}$. In the following, we need to consider every column of matrix $U$ separately, hence, we denote $U=\left[u_{1},\ldots,u_{d}\right],u_{i}\in\mathbb{R}^{\tilde{d}}$. Since, the matrix $U$ is constructed by orthogonal property and its construction is close to $\left\{a^{1},\ldots,a^{t}\right\}_{t\leq d}$. Hence, we denote the Gram-Schmidt Orthogonalization of $\left\{a^{1},\ldots,a^{t}\right\}_{t\leq d}$ as $\left\{v_{1},\ldots,v_{t}\right\}_{t\leq d}$. Then for any $t\leq d$, condition $\operatorname{Span}\left\{v_{1},\ldots,v_{t}\right\}=\operatorname{Span}\left\{a^{1},\ldots,a^{t}\right\}$ holds. Furthermore, for better description of zero-respecting algorithm, we denote $S_{t}=\begin{cases}\cup_{s<t}\operatorname{supp}\left\{\nabla f\left(z^{s}\right)\right\}&t\geq 1\\ \varnothing&t=0\end{cases}$.
Based on these notations, our requirement that the first iterates in sequence $\mathcal{Z}_{\mathcal{A}}\left[f\right]$ and $U^{\top}\mathcal{A}\left[f_{U}\right]$ are identical is equivalent to
$z^{t}=U^{\top}a^{t}\ \text{and}\ \operatorname{supp}\left\{z^{t}\right\}\subseteq S_{t}\ ,\forall t\leq d\ .$ (A.2)
In the following, we begin to prove the equivalent claim (A.2). We explicitly construct algorithm $\mathcal{Z}_{\mathcal{A}}$ and matrix $U$ to meet the requirement in (A.2). The basic idea is that for any subsequence $\left\{a^{t}\right\}_{t\leq d}$, there exists a matrix set $\mathcal{U}\left(t\right)$ and algorithm class $\mathbb{A}\left(t\right)$ such that every matrix $U\in\mathcal{U}\left(t\right)$ and every algorithm $\mathcal{Z}_{\mathcal{A}}\in\mathbb{A}\left(t\right)$ satisfies our requirement. Furthermore, we can show that the intersection of these sets are nonempty, hence, we can always choose the element in the intersection to meet our requirement in (A.2). For the base case, we define the matrix set and function class set as
$\begin{split}\mathcal{U}\left(1\right):=&\left\{U\in\mathbb{R}^{d^{\prime}\times d}:U^{\top}U=I_{d},\left\langle u_{i},v_{1}\right\rangle=0,\forall i\in\left[d\right]\right\}\\ \mathbb{A}\left(1\right):=&\left\{\mathcal{Z}:\mathcal{Z}\left(\varnothing\right)=0\right\}\end{split}$ (A.3)
And the set $\mathcal{U}\left(t\right)$ and $\mathbb{A}\left(t\right)$ is constructed by recursive definition:
$\begin{split}&U\in\mathcal{U}\left(t\right)=\left\{U\in\mathbb{R}^{d\times d}:U^{\top}U=I_{d},\left\langle u_{i},v_{j}\right\rangle=0,\forall i\in\left[d\right]\backslash S_{t},\forall j\leq t\right\}\cap\mathcal{U}\left(t-1\right)\\ &\mathcal{Z}_{\mathcal{A}}\in\mathbb{A}\left(t\right)=\left\{\mathcal{Z}:\begin{array}[]{c}\mathcal{Z}\left(\nabla f\left(z^{1}\right),\ldots,\nabla f\left(z^{t-1}\right)\right)=U^{\top}\mathcal{A}\left(U\nabla f\left(z^{1}\right),\ldots,U\nabla f\left(z^{t-1}\right)\right),\\ U\in\mathcal{U}\left(t\right),\mathcal{Z}\in\mathbb{A}\left(t-1\right)\end{array}\right\},\forall t\geq 2\end{split}$ (A.4)
Before proceeding, we first give more detailed illustration about the construction of the two set $\mathcal{U}\left(t\right)$ and $\mathbb{A}\left(t\right)$. The construction of $\mathbb{A}\left(t\right)$ is to meet the first requirement of (A.2). To keep the relation between that $z^{t}=U^{\top}a^{t}$ holds for any $t$, the algorithm $\mathcal{Z}_{\mathcal{A}}$ must be use the information of initial algorithm $\mathcal{A}$ and try to emulated the output sequence that generated by $\mathcal{A}$. Hence, we select the direct mapping that $\mathcal{Z}_{\mathcal{A}}\left(\nabla\right)=U^{\top}\mathcal{A}\left(U\nabla\right)$. Here we use notation $\nabla$ to denote the information of gradient. The construction of $\mathcal{U}\left(t\right)$ is aim to meet the second requirement of (A.2). Once $z^{t}=U^{\top}a^{t}$ holds, then $\operatorname{supp}\left\{z^{t}\right\}\subseteq S_{t}$ holds only depends on $u_{i},\forall i\in\left[d\right]\backslash S_{t}$ is located in the complementary of $\operatorname{Span}\left\{v_{1},\ldots,v_{t}\right\}$.
Next, we show that every matrix $U\in\mathcal{U}\left(t\right)$ and every algorithm $\mathcal{Z}_{\mathcal{A}}\in\mathbb{A}\left(t\right)$ satisfies our requirement (A.2) with $t$. Let us prove the argument (A.2) with the inductive. Since $a^{1}$ is generated by algorithm $\mathcal{A}$ from scratch, i.e., $a^{1}=\mathcal{A}\left(\varnothing\right)$. Hence, $a^{1}\in\mathbb{R}^{d}$ is an arbitrary (but deterministic) vector. We thus satisfy (A.2) by selecting $U\in\mathcal{U}\left(1\right)$ and $z^{1}=0$, i.e., $0=\mathcal{Z}_{\mathcal{A}}\left(\varnothing\right)$. Obviously, $\mathcal{U}\left(1\right)\neq\varnothing$. Hence, at iteration $t=1$, we can find proper matrix $U$ and $\mathcal{Z}_{\mathcal{A}}$ satisfy (A.2). Suppose construction of $\mathcal{Z}_{\mathcal{A}}\in\mathbb{A}\left(t\right)$ and $U\in\mathcal{U}\left(t\right)$ in (A.4) such that (A.2) holds for any $s<t$, i.e., $z^{s}=U^{\top}a^{s},$ $\operatorname{supp}\left\{z^{s}\right\}\subseteq S_{s}$ and $\mathcal{U}\left(s\right)\neq\varnothing$ for all $s<t\leq T_{0}$. Now, we show that $z^{t}=U^{\top}a^{t}$ and $\operatorname{supp}\left\{z^{t}\right\}\subseteq S_{t}$ hold.
$z^{t}=U^{\top}a^{t}$”: By the induction property and definition of $a^{t}$, $z^{t}$ and $f_{U}\left(a^{t}\right)$, we have that $\nabla f_{U}\left(a^{s}\right)=U\nabla f\left(U^{\top}a^{s}\right)=U\nabla f\left(z^{s}\right)$ for any $s<t$ and $U\in\mathcal{U}\left(t\right)$. Hence,
$z^{t}=\mathcal{Z}_{\mathcal{A}}\left(\nabla f\left(z^{1}\right),\ldots,\nabla f\left(z^{t-1}\right)\right)=U^{\top}\mathcal{A}\left(U\nabla f\left(z^{1}\right),\ldots,U\nabla f\left(z^{t-1}\right)\right)=U^{\top}a^{t}$ (A.5)
holds for any $\mathcal{Z}_{\mathcal{A}}\in\mathbb{A}\left(t\right)$.
$\operatorname{supp}\left\{z^{t}\right\}\subseteq S_{t}$”: Since $z^{t}=U^{\top}a^{t}$, we have $\operatorname{supp}\left\{z^{t}\right\}=\operatorname{supp}\left\{U^{\top}a^{t}\right\}$. To satisfy the support condition that $\operatorname{supp}\left\{z^{t}\right\}\subseteq S_{t}$, we require $\left\langle u_{i},a^{t}\right\rangle=0$ for every $i\notin S_{t}$. Since $a^{t}\in\operatorname{Span}\left\{v_{1},\ldots,v_{t}\right\}$, hence, we need $\left\langle u_{i},v_{j}\right\rangle=0$ for every $i\notin S_{t}$ and $j\leq t\leq d$. Note that $d^{\prime}-\dim\left(\left\{v_{1},\ldots,v_{t}\right\}\right)\geq d\geq|\left[d\right]\backslash S_{t}|$. Hence, we can find proper vectors $u_{i}$ satisfies condition.
Now, we have proven that there exists an algorithm $\mathcal{Z}_{\mathcal{A}}\in\mathbb{A}_{\text{zr}}$ and matrix $U\in\mathbb{R}^{d\times d}$ satisfying the first claim of Lemma A.1. In the following, we show that the algorithm $\mathcal{Z}_{\mathcal{A}}$ and matrix $U$ has property (A.1) for any $\varepsilon$. If $T_{\varepsilon}\left(\mathcal{A},f_{U}\right)>d$, we are done. Otherwise $T_{\varepsilon}\left(\mathcal{A},f_{U}\right)\leq d$ and we have
$\begin{split}&T_{\varepsilon}\left(\mathcal{A},f_{U}\right):=T_{\varepsilon}\left(\mathcal{A}\left[f_{U}\right],f_{U}\right)\\ &=\inf\left\{t\in\mathbb{N}:f_{U}\left(a^{t}\right)-f_{U}\left(a^{*}\right)\leq\varepsilon\right\}\\ &\overset{(a)}{=}\inf\left\{t\in\mathbb{N}:f\left(U^{\top}a^{t}\right)-f^{*}\leq\varepsilon\right\}=T_{\varepsilon}\left(U^{\top}\mathcal{A}\left[f_{U}\right],f\right)\\ &=T_{\varepsilon}\left(U^{\top}\mathcal{A}\left[f_{U}\right],f\right)\overset{(b)}{=}T_{\varepsilon}\left(\mathcal{Z}_{\mathcal{A}},f\right)\end{split}$ (A.6)
[…]
With the help of Lemma A.1, we are ready to prove Proposition 7.1.
Proof of Proposition 7.1.
For any $\mathcal{A}\in\mathbb{A}_{\text{det}}$ and the value $d$, we use Lemma A.1 to construct $\mathcal{Z}_{\mathcal{A}}\in\mathbb{A}_{\text{zr}}$ such that
$T_{\varepsilon}\left(\mathcal{A},f_{U\left(f,\mathcal{A}\right)}\right)\geq\min\left\{d,T_{\varepsilon}\left(\mathcal{Z}_{\mathcal{A}},f\right)\right\}$ (A.7)
for every $f\in\mathcal{F}^{d}$ and some orthogonal matrix $U\left(f,\mathcal{A}\right)\in\mathbb{R}^{\tilde{d}\times d}$ that depends on $f$ and $\mathcal{A}$. Consequently, denote
$\mathcal{F}_{d}^{\tilde{d}}=\left\{f_{U\left(f,\mathcal{A}\right)}:f_{U\left(f,\mathcal{A}\right)}\in\mathcal{F}^{\tilde{d}},f_{U\left(f,\mathcal{A}\right)}=f\left(U\left(f,\mathcal{A}\right)^{\top}x\right),U\left(f,\mathcal{A}\right)\in\mathbb{R}^{\tilde{d}\times d},U\left(f,\mathcal{A}\right)^{\top}U\left(f,\mathcal{A}\right)=I_{d}\right\}$ (A.8)
as a subset of $\mathcal{F}^{\tilde{d}}$, which is parameterized by function $f$ and algorithm $\mathcal{A}$, then we have
$\begin{split}&\mathcal{T}_{\varepsilon}\left(\mathbb{A}_{\text{det}},\mathcal{F}^{\tilde{d}}\right)\\ &=\inf_{\mathcal{A}\in\mathbb{A}_{\text{det}}}\sup_{\tilde{f}\in\mathcal{F}^{\tilde{d}}}T_{\varepsilon}\left(\mathcal{A},\tilde{f}\right)\overset{(a)}{\geq}\inf_{\mathcal{A}\in\mathbb{A}_{\text{det}}}\sup_{f_{U}\in\mathcal{F}_{d}^{\tilde{d}}\subseteq\mathcal{F}^{\tilde{d}}}T_{\varepsilon}\left(\mathcal{A},f_{U}\right)\overset{(b)}{\geq}\inf_{\mathcal{A}\in\mathbb{A}_{\text{det}}}\sup_{f_{U}\in\mathcal{F}_{d}^{\tilde{d}}\subseteq\mathcal{F}^{\tilde{d}}}\min\left\{d,T_{\varepsilon}\left(\mathcal{Z}_{\mathcal{A}},f\right)\right\}\\ &\overset{(c)}{=}\min\left\{d,\inf_{\mathcal{A}\in\mathbb{A}_{\text{det}}}\sup_{f\in\mathcal{F}^{d}}T_{\varepsilon}\left(\mathcal{Z}_{\mathcal{A}},f\right)\right\}\overset{(d)}{\geq}\min\left\{d,\inf_{\mathcal{Z}\in\mathbb{A}_{\text{zr}}}\sup_{f\in\mathcal{F}^{d}}T_{\varepsilon}\left(\mathcal{Z},f\right)\right\}=\min\left\{d,\mathcal{T}_{\varepsilon}\left(\mathbb{A}_{\text{zr}},\mathcal{F}^{d}\right)\right\}\end{split}\ ,$ (A.9)
where $(a)$ holds since the feasible region $\mathcal{F}_{d}^{\tilde{d}}\subseteq\mathcal{F}^{\tilde{d}}$ is a subset of $\mathcal{F}^{\tilde{d}}$ (parameterized by matrix $U$ and function $f$), $(b)$ follows from (A.7), $(c)$ holds by replacing the optimizer from $f_{U}$ to $f$ and $(d)$ is due to $\mathcal{A}\in\mathbb{A}_{\det}$ then the feasible region of $\mathcal{Z}_{\mathcal{A}}$ is a subset of $\mathbb{A}_{\text{zr}}$. As we chose $d$ satisfying $\mathcal{T}_{\varepsilon}\left(\mathbb{A}_{\text{zr}},\mathcal{F}^{d}\right)<d$, hence, inequality (A.9) implies that $\mathcal{T}_{\varepsilon}\left(\mathbb{A}_{\text{det}},\mathcal{F}^{\tilde{d}}\right)\geq\mathcal{T}_{\varepsilon}\left(\mathbb{A}_{\text{zr}},\mathcal{F}^{d}\right)$. ∎∎
Appendix B How to optimize the subproblem
The bundle-level subproblem can be written as
$\min_{x\in\mathbb{R}^{n}}\frac{1}{2}\left\|x-\bar{y}\right\|^{2}\ ,\ \operatornamewithlimits{s.t.}Ax\leq b\ ,$ (B.1)
where each row of $A$ together with the corresponding element of $b$ defines a cut constraint. The quality of our certificate depends critically on the accuracy of this subproblem. Many existing solvers do not meet our accuracy requirement, and the problem often becomes harder as the procedure progresses, frequently due to strong linear dependence among the rows of $A$. It is therefore important to provide an effective solution method.
An important observation is that the dual of (B.1) is
$\min_{\lambda\geq 0}\ \ \frac{1}{2}\lambda^{\top}(AA^{\top})\lambda+\lambda^{\top}\left(b-A\bar{y}\right)\ ,$ (B.2)
which is a nonnegative quadratic program. Furthermore, if there exists $\tilde{b}$ such that $A\tilde{b}=b$, then the problem reduces to
$\min_{\lambda\geq 0}\ \ \frac{1}{2}\left\|A^{\top}\lambda-(\bar{y}-\tilde{b})\right\|^{2}\ ,$ (B.3)
which is a classical nonnegative least squares problem. There is extensive literature on this topic, for example [46, 7, 36]. In addition, an open source Julia package implements related algorithms, see NonNegLeastSquares.jl. In our setting the number of cuts is typically 30 to 50, so the number of rows of $A$ is less than the number of columns (that is, the dimension of the variable), and a vector $\tilde{b}$ satisfying $A\tilde{b}=b$ exists. On the other hand, when such a vector does not exist, the nonnegative quadratic program can still be solved by a modified algorithm for nonnegative least squares with high accuracy. After solving the dual to obtain $\lambda^{*}$, we recover $x^{*}=\bar{y}-A^{\top}\lambda^{*}$ and evaluate the constraint violation $\left\|\left[Ax^{*}-b\right]_{+}\right\|$. If this quantity is below a prescribed tolerance, we accept $x^{*}$; otherwise, we declare the subproblem infeasible.